### v1

# Scattering of the defocusing Calogero–Moser derivative nonlinear Schrödinger equation

Abstract
We consider the defocusing Calogero–Moser derivative nonlinear Schrödinger equation (CM-DNLS)
$i\partial_{t}u+\partial_{x}^{2}u-2\Pi D\left(|u|^{2}\right)u=0,\quad(t,x)\in\mathbb{R}\times\mathbb{R}.$
Using the Gérard-type explicit formula, we prove the scattering result of solutions to this equation with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. We also characterize the scattering using the distorted Fourier transform associated with the Lax operator. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases.
Keywords Calogero–Moser derivative nonlinear Schrödinger equation, Gérard-type explicit formula, Scattering, Distorted Fourier transform
Acknowledgment The author is currently a postdoctoral researcher at the Department of Mathematics and Computer Science, University of Basel. Part of this work was carried out during his PhD studies at the Laboratoire de Mathématiques d’Orsay (UMR 8628), Université Paris-Saclay. The author is deeply grateful to Patrick Gérard for his suggestions throughout this work.
1 Introduction
We consider the following defocusing Calogero–Moser derivative nonlinear Schrödinger equation (CM-DNLS)
$i\partial_{t}u+\partial_{x}^{2}u-2\Pi D\left(|u|^{2}\right)u=0,\quad(t,x)\in\mathbb{R}\times\mathbb{R}.$ (1.1)
Here $\Pi$ denote the Riesz-Szegő orthogonal projector from $L^{2}(\mathbb{R})$ onto
$L_{+}^{2}(\mathbb{R}):=\left\{f\in L^{2}(\mathbb{R}):\operatorname{supp}(\widehat{f})\subset[0,+\infty)\right\},$
and it is given by
$\Pi(f)(x):=\frac{1}{2\pi}\int_{0}^{\infty}e^{i\xi x}\widehat{f}(\xi)d\xi,\quad\forall f\in L^{2}(\mathbb{R}).$ (1.2)
The defocusing Calogero–Moser DNLS equation was first introduced in [22], where it was called the intermediate nonlinear Schrödinger equation. Later, the focusing Calogero–Moser derivative nonlinear Schrödinger equation appeared in [1], where it was obtained as a formal continuum limit of the classical Calogero–Moser systems [20, 21]. In [11], Gérard and Lenzmann studied the following focusing (CM-DNLS),
$i\partial_{t}u+\partial_{x}^{2}u+2\Pi D\left(|u|^{2}\right)u=0,\quad(t,x)\in\mathbb{R}\times\mathbb{R}.$ (1.3)
They proved that if $u\in H_{+}^{s}(\mathbb{R}):=H^{s}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ with $s$ large enough, then equation (1.3) admits a Lax pair structure and has infinitely many conservation laws. Using these conservation laws, they established global well-posedness of (1.3) in $H_{+}^{s}(\mathbb{R})$ for all $s>\tfrac{1}{2}$ when the initial data has small mass ($\|u_{0}\|_{L^{2}}^{2}<2\pi$). Later, Killip, Laurens, and Vişan combined the explicit formulae of (1.1) and (1.3) with methods based on commuting flows to show global well-posedness of both equations in $H_{+}^{s}(\mathbb{R})$ for $s\geq 0$, under the assumption of small mass initial data in the focusing case [17]. Hogan and Kowalski showed that for every $\varepsilon>0$, one can find initial data $u_{0}\in H_{+}^{\infty}$ with mass $2\pi+\varepsilon$ such that the corresponding maximal lifespan solution $u:(T_{-},T_{+})\times\mathbb{R}\to\mathbb{C}$ satisfies $\lim_{t\to T_{\pm}}\|u(t)\|_{H^{s}}=\infty$ for all $s>0$ [16]. K. Kim, T. Kim and Kwon constructed finite-time blow-up solutions in $\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ to (1.3) with mass arbitrarily close to $2\pi$ [18]. More recently, the author showed global well-posedness of (1.1) with a non-vanishing condition at infinity and also established Gérard-type explicit formulae for the corresponding solutions [9].
In [10], Frank and Read proposed an inverse scattering transform for the focusing (CM-DNLS) (1.3) using the distorted Fourier transform. Furthermore, Badreddine characterized the zero dispersion limit of solutions to (1.1) and (1.3) with initial data $u_{0}\in L_{+}^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ through Gérard-type explicit formulae, and identified this limit in terms of the branches of the multivalued solution to the inviscid Burgers–Hopf equation [5]. She also established global well-posedness of (1.1) and (1.3) on the circle in $H_{+}^{s}(\mathbb{T})$ for any $s\geq 0$, under the assumption of small mass initial data in the focusing case [3], and further classified the traveling wave solutions of (1.1) and (1.3) on the circle [4]. Notice that a new numerical scheme based on explicit formulae was recently developped for (1.1) and (1.3) on the circle [2].
In [19], Kim and Kwon established the soliton resolution results for both finite-time blow-up solutions and global-in-time solutions to (1.3). More precisely, they proved scattering of solutions to (1.3) in $H^{1,1}(\mathbb{R}):=\{u\in H^{1}(\mathbb{R}):xu\in L^{2}(\mathbb{R})\}$ with small mass initial data ($\|u_{0}\|_{L^{2}}^{2}<2\pi$). Following their approach, it is expected that solutions to (1.1) in $H^{1,1}(\mathbb{R})$ also scatter as time tends to infinity.
In this paper, we applied the Gérard-type explicit formula to prove the scattering result for solutions to the defocusing (CM-DNLS) (1.1) with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. Our analysis also relies on the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$, and we can characterize the scattering term using this distorted Fourier transform. One can see also [10, Section 4] for the analogous result on the distorted Fourier transform corresponding to the focusing (CM-DNLS) (1.3).
1.1 The Lax pair
In this paper, we denote by $L_{+}^{2}(\mathbb{R})$ the Hardy space corresponding to $L^{2}(\mathbb{R})$ functions having a Fourier transform supported in the domain $\xi\geq 0$. Recall that the space $L_{+}^{2}(\mathbb{R})$ identifies to holomorphic functions on the upper-half plane $\mathbb{C}_{+}:=\{z\in\mathbb{C}:\operatorname{Im}(z)>0\}$ such that
$\sup_{y>0}\int_{\mathbb{R}}|f(x+iy)|^{2}dx<+\infty.$
The Riesz-Szegő projector $\Pi$ is the orthogonal projector from $L^{2}(\mathbb{R})$ onto $L_{+}^{2}(\mathbb{R})$. It is given by
$\forall f\in L^{2}(\mathbb{R}),\quad\forall z\in\mathbb{C}_{+},\quad\Pi f(z)=\frac{1}{2\pi}\int_{0}^{\infty}\mathrm{e}^{iz\xi}\widehat{f}(\xi)d\xi=\frac{1}{2i\pi}\int_{\mathbb{R}}\frac{f(y)}{y-z}dy.$ (1.4)
The Toeplitz operator $T_{b}$ on $L^{2}(\mathbb{R})$ associated to a function $b\in L^{\infty}(\mathbb{R})$ is defined by
$T_{b}f:=\Pi(bf),\qquad\forall f\in L_{+}^{2}(\mathbb{R}).$
For $u\in H_{+}^{1}(\mathbb{R})$, the Lax operator $L_{u}$ corresponding to (1.1) is defined by
$\forall f\in\operatorname{Dom}\left(L_{u}\right)=H_{+}^{1}(\mathbb{R}),\qquad L_{u}f:=Df+uT_{\bar{u}}f\text{ with }D:=\frac{1}{i}\frac{d}{dx}.$
Also, we recall the definition of $X^{*}$ in [14],
$\forall f\in\operatorname{Dom}\left(X^{*}\right):=\left\{f\in L_{+}^{2}(\mathbb{R}):\exists\lambda_{f}\in\mathbb{C}:xf+\lambda_{f}\in L^{2}(\mathbb{R})\right\}=\left\{f\in L_{+}^{2}(\mathbb{R}):\widehat{f}\in H^{1}(0,\infty)\right\},$ (1.5)
with
$\widehat{X^{*}f}(\xi):=i\frac{d}{d\xi}[\widehat{f}(\xi)]\mathbf{1}_{\xi>0}.$
Here $X^{*}$ is the adjoint of the operator of multiplication by $x$ on $L_{+}^{2}(\mathbb{R})$, and $\left(\operatorname{Dom}\left(X^{*}\right),-iX^{*}\right)$ is maximally dissipative. Notice that $iX^{*}$ is the infinitesimal generator of the adjoint semi–group of contractions $\left(S(\eta)^{*}\right)_{\eta\geq 0}$ with
$\forall\eta\geq 0,\qquad S(\eta)^{*}=\mathrm{e}^{-i\eta X^{*}}.$
We also notice that
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\left|\widehat{f}\left(0^{+}\right)\right|^{2}=-4\pi\operatorname{Im}\left\langle X^{*}f\mid f\right\rangle\leq 4\pi\|X^{*}f\|_{L^{2}}\|f\|_{L^{2}}.$
Therefore, we can define
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\quad I_{+}(f):=\widehat{f}\left(0^{+}\right).$ (1.6)
Moreover, we can easily deduce that
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\quad X^{*}f=xf+\frac{1}{2i\pi}I_{+}(f).$ (1.7)
1.2 Gérard-type explicit formula
In [17], Killip, Laurens and Vişan followed Gérard’s approach used in [14] to derive an explicit formula for the solution to (1.1) in $L_{+}^{2}(\mathbb{R})$.
Theorem $1.$
([17, Theorem 1.7]) Let $u\in C\left(\mathbb{R},L_{+}^{2}(\mathbb{R})\right)$ be the solution to (1.1) with $u(0)=u_{0}$. Then for any $t\in\mathbb{R}$,
$\forall z\in\mathbb{C}_{+},\quad u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right].$ (1.8)
The Gérard-type explicit formula was first founded by Gérard in [14] for the Benjamin–Ono equation on the line (see also [8] for the extension), and then by several mathematicians for the cubic Szegő equation on the line [13], the Calogero–Moser DNLS equation [17, 9], and the half wave maps equation [12]. The Gérard-type explicit formula has also been used in several studies [13, 15, 17, 8, 5, 6, 7, 12] so far and has shown to be very effective.
Now we introduce the following lemma concerning the Gérard-type explicit formula, which will be useful in the proof of our main theorem later on.
Lemma 1.2.
Given $u_{0}\in H_{+}^{1}(\mathbb{R})$ and $f_{0}\in L_{+}^{2}(\mathbb{R})$, and we define
$f(t,z):=\frac{1}{2i\pi}I_{+}\left(\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}f_{0}\right),\qquad\forall z\in\mathbb{C}_{+},\,\forall t\in\mathbb{R}.$
Then we have
$|f(t,z)|\leq C(z)\|f\|_{L^{2}(\mathbb{R})}\leq C(z)\|f_{0}\|_{L^{2}(\mathbb{R})},\quad\forall z\in\mathbb{C}_{+},\,\forall t\in\mathbb{R},$ (1.9)
where $C(z)$ depends only on $z$.
Proof.
In [8, Lemma 3.1], the author proved an analogous result corresponding to the Benjamin–Ono equation on the line. One can follow the approach in the proof of [8, Lemma 3.1] to show Lemma 1.2. ∎
1.3 Main result
In this paper, we used the Gérard-type explicit formula (1.8) to study the long-time behavior of (1.1). More precisely, we proved the following scattering result for solutions to (1.1) with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u_{0}\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u_{0}\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$.
Theorem 1.3.
Let $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{w\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. Then the solution $u$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
$\lim_{t\rightarrow+\infty}\left\|u(t)-\mathrm{e}^{i\partial_{x}^{2}}v_{0}\right\|_{L^{2}(\mathbb{R})}=0.$ (1.10)
Here $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$, and
$\mathcal{F}_{d}(f)=\widetilde{f},\qquad\forall f\in L_{+}^{2}(\mathbb{R})$
is the distorted Fourier transform assciated with $L_{u_{0}}$ (see Section 2.1 and Section 2.2 for details).
Remark 1.4.
As mentioned in the introduction, Kim and Kwon proved the scattering result for solutions to (1.3) in $H^{1,1}(\mathbb{R})$ [19]. More precisely, their appraoch relies on the pseudo-conformal symmetry, and thus requires $xu_{0}\in L^{2}(\mathbb{R})$. Also, it is expected that we can follow their approach to show the scattering result for (1.1) in $H^{1,1}(\mathbb{R})$. The great progress we made in this paper is to extend the scattering result to more singular chiral initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. One can observe that the (unique) ground state solution $\frac{\sqrt{2}}{x+i}$ of the focusing (CM-DNLS) (1.3) (see [11, Lemma 4.1]) is included in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\frac{1}{4}<\alpha<\frac{1}{2}$ but not in $H^{1,1}(\mathbb{R})$. On the other hand, we can characterize the scattering term using the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$ of (1.1). Furthermore, this is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. See also [6] for the study on the long-time behavior of the Benjamin–Ono equation with rational initial data and [12] for the study on the long-time behavior of the half-wave maps equation with rational initial data.
Remark 1.5.
Following the strategy in this paper, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [10], we can also establish the scattering result for solutions to (1.3) with initial data in $H_{+}^{1,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ for $\alpha>\frac{1}{4}$.
2 Preliminaries
In this section, we study the generalized eigenfunctions and the distorted Fourier transform associated with the Lax operator $L_{u}$ of (1.1). For $u\in H_{+}^{1}(\mathbb{R})$, it is known that $L_{u}$ has no eigenvalue in this case. Under some additional decay assumption on $u\in H_{+}^{1}(\mathbb{R})$, we show the existence of the generalized eigenfunctions in $L_{+}^{\infty}(\mathbb{R})$ and then introduce the corresponding distorted Fourier transform. See also [10, Section 4] for the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3).
2.1 Generalized eigenfunctions
First we show that if $u\in H_{+}^{1}(\mathbb{R})$, the Lax operator $L_{u}$ has no eigenvalue.
Proposition 2.1.
Let $u\in H_{+}^{1}(\mathbb{R})$, then $\operatorname{ker}\left(L_{u}-\lambda\right)\cap H_{+}^{1}(\mathbb{R})=\{0\}$.
Proof.
Let $\psi\in H_{+}^{1}(\mathbb{R})$ be an eigenvector of $L_{u}$. From the proof of [11, Proposition 5.1], we can infer
$\|\psi\|_{L^{2}}^{2}+\frac{1}{2\pi}|\langle\psi,u\rangle|^{2}=0.$
Thus $\psi=0$. ∎
We now give a fundamental result concerning the generalized eigenfunctions of $L_{u}$, which constitute the foundation of the distorted Fourier transform.
Proposition 2.2.
Let $u\in H^{1}_{+}(\mathbb{R})$ be such that $|x|^{\alpha}u\in L^{2}(\mathbb{R})$ for some $\alpha>0$. For every $\lambda>0$, the space
$\mathcal{M}_{u}(\lambda)=\bigl\{\,m\in L^{\infty}_{+}(\mathbb{R}):\;(L_{u}-\lambda)m=0\ \text{in the distributional sense}\bigr\}$
is one–dimensional. Every element $m\in\mathcal{M}_{u}(\lambda)$ satisfies
$\mathrm{e}^{-i\lambda x}m(x).$
has a limit $\ell_{m}(\pm\infty)\text{ as }x\rightarrow\pm\infty$. Furthermore, $\ell_{m}(-\infty)$ characterizes $m$ and
$\left|\ell_{m}(+\infty)\right|=\left|\ell_{m}(-\infty)\right|.$
Proof.
Notice that, if $m\in L^{\infty}(\mathbb{R})$,
$\Pi(\bar{u}m)\in L_{+}^{2}(\mathbb{R}),$
so that $uT_{\bar{u}}m\in L^{1}(\mathbb{R})$. Equation $(L_{u}-\lambda)m=0$ is equivalent to
$D\left(\mathrm{e}^{-i\lambda x}m(x)\right)=-\mathrm{e}^{-i\lambda x}u(x)\Pi(\bar{u}m)(x),$
which implies that
$D\left(\mathrm{e}^{-i\lambda x}m(x)\right)\in L^{1}(\mathbb{R}),$
hence $\mathrm{e}^{-i\lambda x}m(x)$ has a limit $\ell_{m}(\pm\infty)$ as $x\rightarrow\pm\infty$. Furthermore,
$i\left(\left|\ell_{m}(+\infty)\right|^{2}-\left|\ell_{m}(-\infty)\right|^{2}\right)$ $=\lim_{R\rightarrow+\infty}\int_{-R}^{R}[m(x)\overline{Dm(x)}-Dm(x)\overline{m(x)}]dx$
$=\int_{-\infty}^{\infty}\left[-m(x)\overline{uT_{\bar{u}}m(x)}+uT_{\bar{u}}m(x)\overline{m(x)}\right]dx$
$=-\left\|T_{\bar{u}}m\right\|^{2}+\left\|T_{\bar{u}}m\right\|^{2}$
$=0.$
This proves $\left|\ell_{m}(+\infty)\right|=\left|\ell_{m}(-\infty)\right|$. It remains to prove that the space of such functions $m$ is one dimensional.
Given $\ell\in\mathbb{C}$, the problem $\left(L_{u}-\lambda\right)m=0,\ell_{m}(-\infty)=\ell$ is equivalent to
$m=\mathrm{e}^{i\lambda x}a,\quad a-K_{\lambda}(a)=\ell,$ (2.1)
where $K_{\lambda}:L^{\infty}(\mathbb{R})\rightarrow L^{\infty}(\mathbb{R})$ is defined as
$K_{\lambda}(a)(x):=-i\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}a\right)(y)dy.$
We claim that $1-K_{\lambda}:L^{\infty}\rightarrow L^{\infty}$ is bijective. We shall use the Fredholm alternative by showing that $K_{\lambda}$ is compact and $1-K_{\lambda}$ is injective.
Lemma 2.3.
The operator $K_{\lambda}:L^{\infty}(\mathbb{R})\rightarrow L^{\infty}(\mathbb{R})$ is compact.
Proof.
If $\|a\|_{L^{\infty}}\leq 1$, $\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}a\right)$ belongs to a bounded subset of $L^{1}$. Furthermore,
$\left|K_{\lambda}(a)(x)-K_{\lambda}(a)\left(x^{\prime}\right)\right|\leq\|u\|_{L^{2}}\left(\int_{x}^{x^{\prime}}|u(y)|^{2}dy\right)^{\frac{1}{2}},\qquad x\leq x^{\prime}$
$\left|K_{\lambda}(a)(x)\right|\leq\|u\|_{L^{2}}\left(\int_{-\infty}^{x}|u(x)|^{2}dx\right)^{\frac{1}{2}}$
$\left|K_{\lambda}(a)(x)-K_{\lambda}(a)(+\infty)\right|\leq\|u\|_{L^{2}}\left(\int_{x}^{+\infty}|u(x)|^{2}dx\right)^{\frac{1}{2}}$
For all these reasons, the set
$\left\{K_{\lambda}(a),\|a\|_{L^{\infty}}\leq 1\right\}$
is relatively compact in $L^{\infty}$. ∎
Lemma 2.4.
The operator $\operatorname{Id}-K_{\lambda}:L^{\infty}\rightarrow L^{\infty}$ is injective.
Proof.
Let $f\in L^{\infty}$ such that $f=K_{\lambda}(f)$. Let us prove that $h\in L^{2}(\mathbb{R})$. First, we observe that
$m(x):=\mathrm{e}^{i\lambda x}f(x)$
satisfies $\left(L_{u}-\lambda\right)m=0$ and $\ell_{m}(-\infty)=0$, therefore $\ell_{m}(+\infty)=0$, or equivalently
$\int_{-\infty}^{+\infty}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)dy=0,$
so that
$f(x)=\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)dy=-\int_{x}^{+\infty}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)dy.$
Let $\alpha>0$ such that $|x|^{\alpha}u\in L^{2}$. Then
$g(x):=\bar{u}(x)\mathrm{e}^{i\lambda x}f(x)$
satisfies $\left(1+|x|^{\alpha}\right)g\in L^{2}(\mathbb{R})$, which is equivalent to $\widehat{g}\in H^{\alpha}(\mathbb{R})$. By Hardy’s inequality, $\mathbf{1}_{\mathbb{R}_{+}}\widehat{g}\in H^{\alpha}(\mathbb{R})$, or $\left(1+|x|^{\alpha}\right)\Pi g\in L^{2}(\mathbb{R})$. This implies
$|y|^{2\alpha}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)\in L^{1}(\mathbb{R}),$
and consequently,
$|x|^{2\alpha}|f(x)|\in L^{\infty}(\mathbb{R}).$
If $\alpha>\frac{1}{4}$, this implies $h\in L^{2}(\mathbb{R})$. Otherwise, we plug this information in the above argument. More generally, if $|x|^{\beta}f\in L^{\infty}$ for some $\beta\geq 0$, we have $\left(1+|x|^{\alpha+\beta}\right)f\in L^{2}(\mathbb{R})$, hence, again by Hardy’s inequality,
$\left(1+|x|^{\gamma}\right)g\in L^{2}(\mathbb{R}),$
with $\gamma=\alpha+\beta$ if $\alpha+\beta<\frac{1}{2}$, and $\gamma=\frac{1}{2}-\varepsilon$ for every $\varepsilon>0$ if $\alpha+\beta\geq\frac{1}{2}$. Consequently,
$|x|^{\alpha+\gamma}f\in L^{\infty}$
which leads to $f\in L^{2}$ after a finite number of iterations. Since
$\mathrm{e}^{i\lambda x}f(x)=\int_{-\infty}^{x}\mathrm{e}^{i\lambda(x-y)}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)dy,$
we infer that $m\in L_{+}^{2}\text{, and finally }m\in H_{+}^{1}$. From the assumption on the kernel of $L_{u}-\lambda$, we conclude $m=0$. ∎
Now we can conclude the proof of Proposition 2.2. By the Fredholm alternative, the operator $\operatorname{Id}-K_{\lambda}:L^{\infty}(\mathbb{R})\rightarrow L^{\infty}(\mathbb{R})$ is bejective. Coming back to (2.1), the linear form
$m\in\mathscr{M}_{u}(\lambda)\mapsto\ell_{m}(-\infty)\in\mathbb{C}$
is bijective. Therefore $\mathscr{M}_{u}(\lambda)$ is one dimensional. ∎
We can also conclude the limit of $m(x,\lambda)$ as $\lambda\to 0_{+}$, which is useful in the latter proof.
Lemma 2.5.
Let $\ell_{m}(-\infty)=1$. For any $x\in\mathbb{R}$, we have
$\lim_{\lambda\to 0_{+}}m(x,\lambda)=1.$
Proof.
Let $\lambda=0$ in (2.1), then following the previous operations, we can also prove that the solution to (2.1) with $\lambda=0$ exists and is unique. It is not hard to see that $m=a=1$ is the solution to (2.1) with $\lambda=0$ and $\ell=1$. Now we define $K_{0}:=K_{\lambda}|_{\lambda=0}$, and we can easily infer that
$\forall f\in L^{\infty}(\mathbb{R}),\qquad\lim_{\lambda\to 0_{+}}\|(K_{\lambda}-K_{0})f\|_{L^{\infty}}=0.$
Then from the second resolvent identity, we can deduce that
$\left|m(x,\lambda)-1\right|\leq\left|(\operatorname{Id}-K_{\lambda})^{-1}1-(\operatorname{Id}-K_{0})^{-1}1)\right|\rightarrow 0\text{ as }\lambda\to 0_{+}.$
2.2 The distorted Fourier transform
From Proposition 2.2, for $u\in H_{+}^{1}(\mathbb{R}),|x|^{\alpha}u\in L^{2}(\mathbb{R})$ with some $\alpha>0$, there exists a unique $m_{-}\in\mathscr{M}_{u}(\lambda)$ such that $\ell_{m_{-}}(-\infty)=1$. Given $f\in L_{+}^{1}(\mathbb{R})$, we can define
$\mathcal{F}_{d}f(\lambda):=\widetilde{f}(\lambda):=\int_{\mathbb{R}}f(x)\overline{m_{-}(x,\lambda)}dx,\lambda\geq 0.$
If moreover $f\in L^{2}(\mathbb{R})$, then $\widetilde{f}\in L^{2}([0,+\infty))$. Denote by $\mathcal{H}_{c}\left(L_{u}\right)$ the continuous space of $L_{u}$, namely the space of those $f\in L_{+}^{2}(\mathbb{R})$ having a spectral measure with no pure point part. Denote by $P_{c}$ the orthogonal projector from $L_{+}^{2}$ to $\mathcal{H}_{c}\left(L_{u}\right)$.
Proposition 2.6.
The operator $f\mapsto\widetilde{f}$ extends to $L_{+}^{2}(\mathbb{R})$ with the identity
$\int_{\mathbb{R}}\left|f(x)\right|^{2}dx=\int_{\mathbb{R}}\left|P_{c}f(x)\right|^{2}dx=\frac{1}{2\pi}\int_{0}^{\infty}|\widetilde{f}(\lambda)|^{2}d\lambda.$ (2.2)
Proof.
From the spectral theorem and the fact that $L_{u}$ has no eigenvalue on $0,\infty)$, we have, for every $f\in L_{+}^{2}(\mathbb{R})$,
$\|f\|_{L^{2}}^{2}=\left\|P_{c}f\right\|^{2}=\frac{1}{2i\pi}\lim_{\varepsilon\rightarrow 0^{+}}\int_{0}^{\infty}\left\langle\left(L_{u}-\lambda-i\varepsilon\right)^{-1}f-\left(L_{u}-\lambda+i\varepsilon\right)^{-1}f,f\right\rangle d\lambda.$ (2.3)
Set $g_{\pm}^{\varepsilon}(.,\lambda):=\left(L_{u}-\lambda\mp i\varepsilon\right)^{-1}f$. We have
$g_{\pm}^{\varepsilon}=(D-\lambda\mp i\varepsilon)^{-1}\left(Id+uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}\right)^{-1}f.$
The following lemma can be proved along the same lines as the properties of operator $K_{\lambda}$ in the previous section.
Lemma 2.7.
The following properties hold for every $\lambda\geq 0$, uniformly if $\lambda$ belongs to a compact set of $0,\infty)$.
1. $(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{\infty}$ is uniformly bounded and strongly convergent to $(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{\infty}$ as $\varepsilon\rightarrow 0^{+}$.
2. $uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{1}$ is uniformly compact as $\varepsilon\rightarrow 0^{+}$, and is convergent to $uT_{\bar{u}}(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{1}$ as $\varepsilon\rightarrow 0^{+}$.
3. $\left(Id+uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}\right)^{-1}:L^{1}\rightarrow L^{1}$ is strongly convergent to
$\left(Id+uT_{\bar{u}}(D-\lambda\mp i0)^{-1}\right)^{-1}:L^{1}\rightarrow L^{1}$
as $\varepsilon\rightarrow 0^{+}$.
Assume that $f\in L_{+}^{1}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$. From this lemma, we infer that $g_{\pm}^{\varepsilon}$ tends to $g_{\pm}$ in $L^{\infty}$ with
$g_{+}(x,\lambda)=\int_{-\infty}^{x}\mathrm{e}^{i\lambda(x-y)}f(y)dy-\int_{-\infty}^{x}\mathrm{e}^{i\lambda(x-y)}u(y)\Pi_{+}\left(\bar{u}g_{+}(.,\lambda)\right)(y)dy$
$g_{-}(x,\lambda)=-\int_{x}^{+\infty}\mathrm{e}^{i\lambda(x-y)}f(y)dy+\int_{x}^{+\infty}\mathrm{e}^{i\lambda(x-y)}u(y)\Pi_{+}\left(\bar{u}g_{-}(.,\lambda)\right)(y)dy$
In particular, from Proposition 2.2, $g_{+}-g_{-}=-\ell_{g_{-}}(-\infty)m_{-}$, with
$-i\ell_{g_{-}}(-\infty)=\lim_{R\rightarrow+\infty}\int_{-R}^{R}\left(g_{-}\overline{Dm_{-}}-Dg_{-}\bar{m}_{-}\right)dx=-\int_{-\infty}^{\infty}f\bar{m}_{-}dx=-\widetilde{f}(\lambda)$
and finally
$g_{+}(x,\lambda)-g_{-}(x,\lambda)=i\widetilde{f}(\lambda)m_{-}(x,\lambda)$
and, passing to the limit in the right hand side of (2.3) leads to (2.2) for $f\in L_{+}^{1}\cap L_{+}^{2}$. By density, the operator $f\mapsto\widetilde{f}$ extends to $L_{+}^{2}(\mathbb{R})$, with identity (2.2). ∎
Remark 2.8.
Notice that, if $f\in H_{+}^{1}(\mathbb{R})$,
$\widetilde{L_{u}f}(\lambda)=\int_{\mathbb{R}}L_{u}f(x)\overline{m_{-}(x,\lambda)}dx=\int_{\mathbb{R}}f(x)\overline{L_{u}m_{-}(x,\lambda)}dx=\lambda\widetilde{f}(\lambda).$
Applying the spectral theorem to the self-adjoint operator $L_{u}$ restricted to $H_{c}(L_{u})$, combined with Proposition 2.2, we can infer the surjectivity of the distorted Fourier transform.
Corollary 2.9.
The map $f\in L_{+}^{2}(\mathbb{R})\mapsto\widetilde{f}\in L^{2}(0,\infty)$ is onto.
Using the interpolation between $\mathcal{F}_{d}:L_{+}^{1}(\mathbb{R})\to L^{\infty}(0,\infty)$ and $\mathcal{F}_{d}:L_{+}^{2}(\mathbb{R})\to L^{2}(0,\infty)$, we can also extend the distorted Fourier transform $\mathcal{F}_{d}$ to $\mathcal{F}_{d}:L_{+}^{p}\left(\mathbb{R}\right)\rightarrow L^{p^{\prime}}\left(0,\infty\right)$ for $1<p<2$.
Corollary 2.10.
The distorted Fourier transform can be extended uniquely and continuously to
$\mathcal{F}_{d}:L_{+}^{p}\left(\mathbb{R}\right)\rightarrow L^{p^{\prime}}\left(0,\infty\right),\qquad 1\leq p\leq 2,$
where
$\frac{1}{p}+\frac{1}{p^{\prime}}=1.$
As is well known, the classical Fourier transform is a continuous bijection from $\mathcal{S}(\mathbb{R})$ onto itself. Similarly, for the distorted Fourier transform defined above, if we restrict to $u\in\mathcal{S}_{+}(\mathbb{R})$, we obtain an analogous property.
Proposition 2.11.
For $u\in\mathcal{S}_{+}(\mathbb{R})$, the distorted Fourier transform $\mathcal{F}_{d}:\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})\rightarrow\mathcal{S}_{*}:=\{f=\left.g\right|_{(0,\infty)}:g\in\mathcal{S}(\mathbb{R})\text{ and }\operatorname{supp}(g)\subset[0,\infty)\}$ is well-defined and a continuous bijection.
Proof.
Let us recall (2.1) with $\ell=1$,
$m_{-}=\mathrm{e}^{i\lambda x}a,\qquad a-K_{\lambda}(a)=1\text{ with }K_{\lambda}(a)(x):=-i\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}a\right)(y)dy.$
so we have
$a=(Id-K_{\lambda})^{-1}(1).$
Then
$\partial_{x}a(x,\lambda)=-i\mathrm{e}^{-i\lambda x}u(x)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}a\right)(x)\in L_{x,\lambda}^{\infty}.$
From the above expression, it is not hard to see that
$\forall n\in\mathbb{N}_{+},\qquad\partial_{x}^{n}a(x,\lambda)\in L_{x,\lambda}^{\infty}.$
Also,
$\partial_{\lambda}a(x,\lambda)$ $=(Id-K_{\lambda})^{-1}\partial_{\lambda}K_{\lambda}(Id-K_{\lambda})^{-1}(1)$
$=(Id-K_{\lambda})^{-1}\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u[T_{\bar{u}},y]\mathrm{e}^{i\lambda y}a(y,\lambda)dy$
$=\frac{1}{2i\pi}(Id-K_{\lambda})^{-1}\left(\bar{\widetilde{u}}(\lambda)\int_{-\infty}^{x}u(y)\mathrm{e}^{-i\lambda y}dy\right),$
here the last identity comes from
$\forall f\in L^{\infty},\qquad u[T_{\bar{u}},y]f=\frac{u}{2i\pi}\int_{\mathbb{R}}f\bar{u}dx.$
Since
$\left|\lambda\widetilde{u}(\lambda)\right|$ $=\left|\lambda\int_{\mathbb{R}}u(x)\mathrm{e}^{-ix\lambda}\bar{a}(x,\lambda)dx\right|$
$=\left|\int_{\mathbb{R}}u(x)\partial_{x}(\mathrm{e}^{-ix\lambda})\bar{a}(x,\lambda)dx\right|$
$\leq\left|\int_{\mathbb{R}}u^{\prime}(x)\mathrm{e}^{-ix\lambda}\bar{a}(x,\lambda)dx\right|+\left|\int_{\mathbb{R}}u(x)\mathrm{e}^{-ix\lambda}\partial_{x}\bar{a}(x,\lambda)dx\right|$
$\leq C,$
we deduce that
$\|\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-1}.$
Furthermore, it is not hard to conclude
$\forall k\in\mathbb{N}_{+},\qquad\|\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq C_{k}(1+|\lambda|)^{-k}.$
Moreover, since
$\partial_{x}\partial_{\lambda}a=\frac{1}{2i\pi}\widetilde{u}(\lambda)u(x)\mathrm{e}^{-i\lambda x}-i\mathrm{e}^{-i\lambda x}u(x)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}\partial_{\lambda}a\right)(x),$
we have
$\|\partial_{x}\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq\frac{1}{2\pi}\|\widetilde{u}(\lambda)u(x)\|_{L_{x}^{\infty}}+\|\mathrm{e}^{-i\lambda x}u(x)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}\partial_{\lambda}a\right)(x)\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-1}.$
It is not hard to conclude
$\forall n\in\mathbb{N}_{+},\qquad\|\partial_{x}^{n}\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-1}.$
For higher $\lambda$–derivatives, differentiating repeatedly in $\lambda$ produces sums of terms where derivatives fall on $\mathrm{e}^{-i\lambda y},\mathrm{e}^{i\lambda y}$ and on $a$. Each time we gain a primitive like $\int_{-\infty}^{x}y^{k}u(y)e^{-i\lambda y}dy$ (which decays faster than any power of $|\lambda|$ since $u\in\mathcal{S}_{+}(\mathbb{R})$), $\int_{\mathbb{R}}y^{k}\bar{u}(y)e^{i\lambda y}\partial_{\lambda}^{\ell}a(y,\lambda)dy$ or $\partial_{\lambda}^{k}\bar{\widetilde{u}}(\lambda)$. Then we can follow the previous argument and use the recurrence to show that
$\forall n\in\mathbb{N}_{+}\text{ and }\forall m\in\mathbb{N}_{+},\qquad\|\partial_{x}^{n}\partial_{\lambda}^{m}a\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-m}.$
So $a$ is a symbol of order 0, by the classical pseudo-differential operators theory, we can conclude that the distorted Fourier transform $\mathcal{F}_{d}:\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})\rightarrow\mathcal{S}_{*}:=\{f=\left.g\right|_{(0,\infty)}:g\in\mathcal{S}(\mathbb{R})\text{ and }\operatorname{supp}(g)\subset[0,\infty)\}$ is well-defined and a continuous map. Proposition 2.6 implies the injection of $\mathcal{F}_{d}$. Since the distorted Fourier transform is a unitary operator, we can easily infer the surjection of $\mathcal{F}_{d}$ from $\mathcal{S}_{+}(\mathbb{R})$ to $\mathcal{S}_{*}$. ∎
Corollary 2.12.
For $u\in\mathcal{S}_{+}(\mathbb{R})$, the distorted Fourier transform $\mathcal{F}_{d}$ extends uniquely and continuously to the tempered distributions: $\mathcal{F}_{d}:\mathcal{S}_{+}^{\prime}(\mathbb{R})\to\mathcal{S}_{*}^{\prime}(\mathbb{R})$, via
$\langle\mathcal{F}_{d}F,\ \phi\rangle:=2\pi\langle F,\ \mathcal{F}_{d}^{-1}\phi\rangle,\qquad\forall F\in\mathcal{S}^{\prime}_{+}(\mathbb{R}),\phi\in\mathcal{S}_{*}(\mathbb{R}).$
Proof.
By Proposition 2.11 we know that the operator
$\mathcal{F}_{d}f(\lambda)=\int_{\mathbb{R}}e^{-i\lambda x}\overline{a(x,\lambda)}f(x)\,dx$
maps $\mathcal{S}_{+}(\mathbb{R})$ continuously onto $\mathcal{S}_{*}$. Its inverse on $\mathcal{S}_{*}$ is given by
$\mathcal{F}_{d}^{-1}\varphi(x)=\frac{1}{2\pi}\int_{0}^{\infty}e^{i\lambda x}a(x,\lambda)\,\varphi(\lambda)\,d\lambda,$
which is continuous $\mathcal{S}_{*}\to\mathcal{S}_{+}$ and satisfies $\mathcal{F}_{d}^{-1}\circ\mathcal{F}_{d}=\mathrm{Id}_{\mathcal{S}_{+}}$ and $\mathcal{F}_{d}\circ\mathcal{F}_{d}^{-1}=\mathrm{Id}_{\mathcal{S}_{*}}$.
For $F\in\mathcal{S}_{+}^{\prime}(\mathbb{R})$, we can then define the distorted Fourier transform of this distribution by
$\langle\mathcal{F}_{d}F,\ \phi\rangle:=2\pi\langle F,\ \mathcal{F}_{d}^{-1}\phi\rangle,\qquad\forall F\in\mathcal{S}^{\prime}_{+}(\mathbb{R}),\phi\in\mathcal{S}_{*}(\mathbb{R}).$
Here $\mathcal{F}_{d}F\in\mathcal{S}_{*}^{\prime}(\mathbb{R})$. ∎
Now we go to analyse the form of $\partial_{\lambda}a$. We know that, the problem $(L_{u}-\lambda)m_{-}=0,\ell_{m_{-}}(-\infty)=1$ is equivalent to
$m(\lambda,x)=e^{i\lambda x}a(x,\lambda),\qquad a(x,\lambda)=1+K_{\lambda}(a)(x),$
where
$K_{\lambda}(a)(x):=-i\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}(y)\mathrm{e}^{i\lambda y}a(\lambda,y)\right)dy.$
Again, we assume that $u\in\mathcal{S}_{+}(\mathbb{R})$. We have $\partial_{\lambda}a$ is well defined in $L_{+}^{\infty}$ and
$\partial_{\lambda}a=\frac{1}{2i\pi}\bar{\widetilde{u}}(\lambda)\int_{-\infty}^{x}e^{-i\lambda y}u(y)dy+K_{\lambda}(\partial_{\lambda}a).$
Denote $n(x,\lambda):=e^{i\lambda x}\partial_{\lambda}a$, we infer
$L_{u}n-\lambda n=-\frac{1}{2\pi}\bar{\widetilde{u}}(\lambda)u(x).$
Applying the distorted Fourier transform and Proposition 2.2, we can deduce that
$\widetilde{n}(\mu,\lambda)=-\frac{1}{2\pi}p.v.\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)}{\mu-\lambda}+c\delta(\mu-\lambda),$
i.e.,
$n(x,\lambda)=n_{1}(x,\lambda)+c\,m_{-}(x,\lambda),$ (2.4)
where
$\widetilde{n}_{1}(\mu,\lambda):=-\frac{1}{2\pi}p.v.\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)}{\mu-\lambda}$
and $c$ to be determined.
To obtain $c$, we need to give the aysmptotic behavior of $\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$ at $x=$-$\infty$. For this purpose, we establish the following lemma.
Lemma 2.13.
Let $u\in\mathcal{S}_{+}(\mathbb{R})$, then we have
$\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$
Proof.
By the classical Fourier transform theory, we know that
$p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)e^{ix\mu}}{\mu-\lambda}d\mu$
$=p.v.\int_{0}^{\infty}\frac{|\widetilde{u}(\lambda)|^{2}e^{ix\mu}}{\mu-\lambda}d\mu+p.v.\int_{0}^{\infty}\bar{\widetilde{u}}(\lambda)\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}e^{ix\mu}d\mu$
$=e^{i\lambda x}|\widetilde{u}(\lambda)|^{2}\left[-\mathrm{Ci}(\lambda|x|)+i\operatorname{sgn}(x)\left(\frac{\pi}{2}+\mathrm{Si}(\lambda|x|)\right)\right]+p.v.\int_{0}^{\infty}\bar{\widetilde{u}}(\lambda)\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}e^{ix\mu}d\mu,$
where
$\operatorname{Si}(w)=\int_{0}^{w}\frac{\sin t}{t}dt,\qquad\operatorname{Ci}(w)=-\int_{w}^{\infty}\frac{\cos t}{t}dt.$
Since $\widetilde{u}\in\mathcal{S}_{*}$, we know that
$\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}\in L^{1}([0,\infty)),$
so we infer that
$p.v.\int_{0}^{\infty}\bar{\widetilde{u}}(\lambda)\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}e^{ix\mu}d\mu\longrightarrow 0\text{ as }x\rightarrow-\infty.$
Thus we have
$-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)e^{ix\mu}}{\mu-\lambda}d\mu\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$
Now we analyze $\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$. According to (2.1), we have
$m_{-}(x,\mu)=\mathrm{e}^{i\mu x}+\mathrm{e}^{i\mu x}K_{\mu}a(x):=\mathrm{e}^{i\mu x}-i\mathrm{e}^{i\mu x}\int_{-\infty}^{x}\mathrm{e}^{-i\mu y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\mu\cdot}a\right)(y)dy.$
Then
$\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$ $:=-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)e^{ix\mu}}{\mu-\lambda}d\mu-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)\mathrm{e}^{i\mu x}K_{\mu}a(x)}{\mu-\lambda}d\mu$
$:=I_{1}(x,\lambda)+I_{2}(x,\lambda).$
with
$I_{2}(x,\lambda)$
$=-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{2\lambda}\frac{|\widetilde{u}(\lambda)|^{2}\mathrm{e}^{i\mu x}K_{\lambda}a(x)}{\mu-\lambda}d\mu-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{2\lambda}\frac{\bar{\widetilde{u}}(\lambda)\mathrm{e}^{i\mu x}\left(\widetilde{u}(\mu)K_{\mu}a(x)-\widetilde{u}(\lambda)K_{\lambda}a(x)\right)}{\mu-\lambda}d\mu$
$-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{2\lambda}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)\mathrm{e}^{i\mu x}K_{\mu}a(x)}{\mu-\lambda}d\mu$
$:=I_{3}(x,\lambda)+I_{4}(x,\lambda)+I_{5}(x,\lambda).$
We already know that
$I_{1}(x,\lambda)\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$ (2.5)
Since $K_{\lambda}a(x)\rightarrow 0$ as $x\rightarrow$-$\infty$, it is not hard to see that
$I_{3}(x,\lambda)\longrightarrow 0\text{ as }x\rightarrow-\infty.$ (2.6)
Since $u\in\mathcal{S}_{+}(\mathbb{R})$ and $\widetilde{u}\in\mathcal{S}_{*}$, we can easily deduce that for $\mu,\lambda>0$,
$|\widetilde{u}(\mu)K_{\mu}a(x)-\widetilde{u}(\lambda)K_{\lambda}a(x)|\leq C(x)|\mu-\lambda|,$
where $C(x)$ does not depend on $\mu,\lambda$ and $C(x)\rightarrow 0$ as $x\rightarrow$-$\infty$.
Then we infer
$|I_{4}|\leq\frac{2\lambda|\widetilde{u}(\lambda)|}{4\pi^{2}}C(x)\rightarrow 0\text{ as }x\rightarrow-\infty.$ (2.7)
Since $\|K_{\mu}a(x)\|_{L_{\mu}^{\infty}}\rightarrow 0$ as $x\rightarrow$-$\infty$ by the Cauchy–Schwarz inequality, we can easily deduce that
$|I_{5}|\leq C_{\lambda}\|\widetilde{u}\|_{L_{\mu}^{2}}\|K_{\mu}a(x)\|_{L_{\mu}^{\infty}}\rightarrow 0\text{ as }x\rightarrow$-$\infty$ (2.8)
Combining (2.5), (2.6), (2.7) and (2.8), we conclude that
$\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$
The proof is complete. ∎
From Lemma 2.13, we can then easily give the form of $n(x,\lambda)$.
Corollary 2.14.
Let $u\in\mathcal{S}_{+}(\mathbb{R})$, then we have the following expression of $n(x,\lambda)$
$n(x,\lambda)=n_{1}(x,\lambda)-\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}m_{-}(x,\lambda)$ (2.9)
Proof.
Recall the expression (2.4)
$n(x,\lambda)=n_{1}(x,\lambda)+c\,m_{-}(x,\lambda),$
since $n(x,\lambda)\rightarrow 0$ as $x\rightarrow$-$\infty$, from Lemma 2.13, we can infer that
$c=-\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}m_{-}(x,\lambda).$
We keep the assumption $u\in\mathcal{S}_{+}(\mathbb{R})$. Let $\widetilde{f}\in H^{1}(0,\infty)$, $f\in\operatorname{Dom}(X^{*})$ and $f\in H_{+}^{1}(\mathbb{R})$, then
[…]
$=-i\widetilde{f}(0_{+})-i\lim_{\lambda\to 0_{+}}\int_{0}^{\infty}\widetilde{f}(\lambda)\,\partial_{\lambda}\chi_{\varepsilon}(\lambda)m_{-}(x,\lambda)d\lambda-i\lim_{\lambda\to 0_{+}}\int_{0}^{\infty}\widetilde{f}(\lambda)\,\chi_{\varepsilon}(\lambda)\partial_{\lambda}m_{-}(x,\lambda)d\lambda$
$=-i\widetilde{f}(0_{+})-i\int_{0}^{\infty}\widetilde{f}(\lambda)\,\partial_{\lambda}m_{-}(x,\lambda)d\lambda$
$=-i\widetilde{f}(0_{+})+x\int_{0}^{\infty}\widetilde{f}(\lambda)\,m_{-}(x,\lambda)d\lambda-i\int_{0}^{\infty}\widetilde{f}(\lambda)n(x,\lambda)d\lambda.$
Here $\chi_{\varepsilon}(\lambda)=\mathrm{e}^{-\varepsilon\lambda^{2}}$, and the second identity comes from the integral by parts and Lemma 2.5.
From Corollary 2.14, we deduce
[…]
$=-\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)n_{1}(x,\lambda)d\lambda-\frac{1}{8\pi^{2}}\int_{0}^{\infty}\widetilde{f}(\lambda)|\widetilde{u}(\lambda)|^{2}m_{-}(x,\lambda)d\lambda$
$=-\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)n_{1}(x,\lambda)d\lambda-\frac{1}{4\pi}\mathcal{F}_{d}^{-1}(\widetilde{f}(\lambda)|\widetilde{u}(\lambda)|^{2}).$
Let us denote the third term above by $r_{3}(x)$. Then we can easily compute its distorted Fourier transform:
$\widetilde{r}_{3}(\mu)=-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)\widetilde{n_{1}}(\mu,\lambda)d\lambda=\frac{i}{4\pi}\widetilde{u}(\mu)H(\bar{\widetilde{u}}\widetilde{f})(\mu).$
Here $H$ is the Hilbert transform on the positive half-line.
Since $u\in\mathcal{S}_{+}(\mathbb{R})$, then we can observe that
$-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)n_{1}(x,\lambda)d\lambda-\frac{1}{4\pi}\mathcal{F}_{d}^{-1}(\widetilde{f}(\lambda)|\widetilde{u}(\lambda)|^{2})=\frac{1}{4\pi}\mathcal{F}_{d}^{-1}\left(i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{f})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{f}(\lambda)\right)\in L_{+}^{2}(\mathbb{R}).$
Also we have $\mathcal{F}_{d}^{-1}(\partial_{\lambda}\widetilde{f})\in L_{+}^{2}(\mathbb{R})$, therefore
$-\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf\in L_{+}^{2}(\mathbb{R}).$
Then from the definition of $X^{*}$ in (1.5) and the identity (1.7), we can deduce that $\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf=X^{*}f$ and
$\widetilde{f}(0_{+})=\widehat{f}(0_{+}).$ (2.10)
Thus we have
$i\partial_{\lambda}\widetilde{f}=\mathcal{F}_{d}\left(X^{*}f\right)+\frac{1}{4\pi}\left(i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{f})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{f}(\lambda)\right).$ (2.11)
Even in the distorted Fourier transform, the resolvent of $X^{*}f+2tL_{u_{0}}$ is still difficult to compute. However, we can easily compute the resolvent of $i\partial_{\lambda}+2t\lambda$ with the following form
$\widetilde{h}(\lambda):=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}(\eta)d\eta.$
For every $z\in\mathbb{C}_{+}$, it is not hard to see that $\widetilde{h}\in H^{1}(0,\infty),h\in H_{+}^{1}(\mathbb{R})$ and $h\in\operatorname{Dom}(X^{*})$ using the Schur test. Now we justify $I_{+}(h)$. From (2.10), we have
$\frac{1}{2i\pi}I_{+}(h)=\widehat{h}(0_{+})=\widetilde{h}(0_{+})=\frac{1}{2\pi}\int_{0}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}(\eta)d\eta=\mathrm{e}^{it\partial_{x}^{2}}v(z),$ (2.12)
where $\widehat{v}=\widetilde{u}$.
From the density of $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$, we can obtain (2.12) for $u\in L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$.
3 Long-time behavior
Now we can prove the following scattering result for the the defocusing (CM-DNLS) (1.1).
Theorem 3.1.
Let $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. Then the solution $u$ to the defocusing (CM-DNLS) (1.1) satisfies
$\lim_{t\rightarrow+\infty}\left\|u(t)-\mathrm{e}^{i\partial_{x}^{2}}v_{0}\right\|_{L^{2}(\mathbb{R})}=0.$ (3.1)
Here $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$, and
$\mathcal{F}_{d}(f)=\widetilde{f},\qquad\forall f\in L_{+}^{2}(\mathbb{R})$
is the distorted Fourier transform assciated with $L_{u_{0}}$.
Proof.
Since $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R})$, we infer $u_{0}\in L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$. From Corollary 2.10, we know that $\widetilde{u}_{0}\in L^{4}(0,\infty)$ and
$\|\widetilde{u}_{0}\|_{L^{4}(0,\infty)}\leq C\|u\|_{L^{\frac{4}{3}}(\mathbb{R})}.$ (3.2)
The mass conservation law $\|u(t)\|_{L^{2}(\mathbb{R})}=\|u_{0}\|_{L^{2}(\mathbb{R})}$ implies that u(t){u(t)} is uniformly bounded in $L_{+}^{2}(\mathbb{R})$, hence it admits weakly convergent subsequences in $L_{+}^{2}(\mathbb{R})$ as $t\to\infty$. We shall prove that the weak limit is unique.
Let us recall the explicit formula of solution to the defocusing (CM-DNLS) (1.1),
$\forall z\in\mathbb{C}_{+},\quad u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right].$ (3.3)
Now we denote
$\mathcal{F}_{d}^{-1}(i\partial_{\lambda}\widetilde{f}):=Yf,\hskip 18.49988pt\forall\widetilde{f}\in H^{1}(0,\infty).$
We recall (2.12) and we rewrite it as
$h_{0}:=(Y+2tL_{u_{0}}-zId)^{-1}u_{0}$
and
$\frac{1}{2i\pi}I_{+}\left[(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right]=\frac{1}{2i\pi}I_{+}\left[h_{0}\right]=\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z).$
Then we have
$\,u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)$
$=$ $\,\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right]-\frac{1}{2i\pi}I_{+}\left[(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right]$
$=$ $\,\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}(Y-X^{*})(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right].$
The last identity above comes from the second resolvent identity.
Then from Lemma 1.2, (2.11) and (3.2), we infer
$|u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)|$
$\leq\,$ $C(z)\left\|(Y-X^{*})(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right\|_{L^{2}(\mathbb{R})}$
$\leq\,$ $C(z)\left\|\mathcal{F}_{d}\left((Y-X^{*})(Y+2tL_{u_{0}}-zId)^{-1}u_{0}\right)\right\|_{L^{2}(0,\infty)}$
$\leq\,$ $C(z)\left\|i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{h}_{0})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{h}_{0}(\lambda)\right\|_{L^{2}(0,\infty)}$
$\leq\,$ $C(z)\|\widetilde{u_{0}}\|_{L^{4}(0,\infty)}^{2}\|\widetilde{h}_{0}(t)\|_{L_{\lambda}^{\infty}(0,\infty)}$
$\leq\,$ $C(z)\|u_{0}\|_{L^{\frac{4}{3}}(0,\infty)}^{2}\|\widetilde{h}_{0}(t)\|_{L_{\lambda}^{\infty}(0,\infty)},$
where
$\widetilde{h_{0}}(t,\lambda)=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}_{0}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u_{0}}(\eta)d\eta.$
Since $z\in\mathbb{C}_{+}$, Now we apply the Van der Corput lemma and we can conclude that
$\lim_{t\to\infty}\|\widetilde{h}_{0}(t)\|_{L_{\lambda}^{\infty}(0,\infty)}=0$
Thus we deduce
$\forall z\in\mathbb{C}_{+},\qquad u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)\to 0\text{ as }t\to\infty.$
If $u(t_{n})\rightharpoonup w$ in $L_{+}^{2}(\mathbb{R})$ as $t_{n}\to\infty$, then for every $z\in\mathbb{C}_{+}$, we have
$u(t_{n},z)-w(z)\to 0\text{ as }t_{n}\to\infty.$
Therefore we can deduce $w=\mathrm{e}^{it\partial_{x}^{2}}v_{0}$ and the weak limit is unique, i.e.,
$u(t)\rightharpoonup\mathrm{e}^{it\partial_{x}^{2}}v_{0}\text{ in }L_{+}^{2}(\mathbb{R})\text{ as }t\to\infty.$
Since $\|u(t)\|_{L^{2}}=\|u_{0}\|_{L^{2}}=\|\mathrm{e}^{it\partial_{x}^{2}}v_{0}\|_{L^{2}}$, we conclude that this weak limit is exactly the strong limit in $L_{+}^{2}(\mathbb{R})$. The proof is complete. ∎