### v3

# Scattering of the defocusing Calogero–Moser derivative nonlinear Schrödinger equation

Abstract
We consider the defocusing Calogero–Moser derivative nonlinear Schrödinger equation (CM-DNLS)
$i\partial_{t}u+\partial_{x}^{2}u-2\Pi D\left(|u|^{2}\right)u=0,\quad(t,x)\in\mathbb{R}\times\mathbb{R}.$
Using the Gérard-type explicit formula, we prove the scattering result of solutions to this equation with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. We also characterize the scattering term using the distorted Fourier transform associated with the Lax operator. Following our approach developed in this paper, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases.
Keywords Calogero–Moser derivative nonlinear Schrödinger equation, Gérard-type explicit formula, Scattering, Distorted Fourier transform
Acknowledgment The author is currently a postdoctoral researcher at the Department of Mathematics and Computer Science, University of Basel. Part of this work was carried out during his PhD studies at the Laboratoire de Mathématiques d’Orsay (UMR 8628), Université Paris-Saclay. The author is deeply grateful to Patrick Gérard for his insightful comments throughout this work. The author also wishes to thank Thierry Laurens for his helpful suggestions.
1 Introduction
We consider the following defocusing Calogero–Moser derivative nonlinear Schrödinger equation (CM-DNLS)
$i\partial_{t}u+\partial_{x}^{2}u-2\Pi D\left(|u|^{2}\right)u=0,\quad(t,x)\in\mathbb{R}\times\mathbb{R}.$ (1.1)
Here $\Pi$ denote the Riesz-Szegő orthogonal projector from $L^{2}(\mathbb{R})$ onto
$L_{+}^{2}(\mathbb{R}):=\left\{f\in L^{2}(\mathbb{R}):\operatorname{supp}(\widehat{f})\subset[0,+\infty)\right\},$
and it is given by
$\Pi(f)(x):=\frac{1}{2\pi}\int_{0}^{\infty}e^{i\xi x}\widehat{f}(\xi)d\xi,\quad\forall f\in L^{2}(\mathbb{R}).$ (1.2)
The defocusing Calogero–Moser DNLS equation was first introduced in [27], where it was called the intermediate nonlinear Schrödinger equation. Later, the focusing Calogero–Moser derivative nonlinear Schrödinger equation appeared in [1], where it was obtained as a formal continuum limit of the classical Calogero–Moser systems [23, 24]. In [12], Gérard and Lenzmann studied the following focusing (CM-DNLS),
$i\partial_{t}u+\partial_{x}^{2}u+2\Pi D\left(|u|^{2}\right)u=0,\quad(t,x)\in\mathbb{R}\times\mathbb{R}.$ (1.3)
They proved that if $u\in H_{+}^{s}(\mathbb{R}):=H^{s}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ with $s$ large enough, then the equation (1.3) admits a Lax pair structure and has infinitely many conservation laws. Using these conservation laws, they established global well-posedness of (1.3) in $H_{+}^{s}(\mathbb{R})$ for all $s>\tfrac{1}{2}$ when the initial data has small mass ($\|u_{0}\|_{L^{2}}^{2}<2\pi$). Later, Killip, Laurens, and Vişan combined the explicit formulae of (1.1) and (1.3) with methods based on commuting flows to show global well-posedness of both equations in $H_{+}^{s}(\mathbb{R})$ for $s\geq 0$, under the assumption of small mass initial data in the focusing case [19]. Hogan and Kowalski showed that for every $\varepsilon>0$, one can find initial data $u_{0}\in H_{+}^{\infty}$ with mass $2\pi+\varepsilon$ such that the corresponding maximal lifespan solution $u:(T_{-},T_{+})\times\mathbb{R}\to\mathbb{C}$ satisfies $\lim_{t\to T_{\pm}}\|u(t)\|_{H^{s}}=\infty$ for all $s>0$ [17]. K. Kim, T. Kim and Kwon constructed finite-time blow-up solutions in $\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$ to (1.3) with mass arbitrarily close to $2\pi$ [20]. The quantized blow-up dynamics for the focusing (CM-DNLS) (1.3) was studied in [18]. More recently, the author showed global well-posedness of (1.1) with a non-vanishing condition at infinity and also established Gérard-type explicit formulae for the corresponding solutions [10].
In [22], Matsuno studied the Cauchy problem for the defocusing (CM-DNLS) (1.1) with a non-zero constant condition at infinity, developing an inverse scattering transform formulation. In [11], Frank and Read proposed an inverse scattering transform for the focusing (CM-DNLS) (1.3) using the distorted Fourier transform.
Furthermore, Badreddine characterized the zero dispersion limit of solutions to (1.1) and (1.3) with initial data $u_{0}\in L_{+}^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ through Gérard-type explicit formulae, and identified this limit in terms of the branches of the multivalued solution to the inviscid Burgers–Hopf equation [5]. She also established global well-posedness of (1.1) and (1.3) on the circle in $H_{+}^{s}(\mathbb{T})$ for any $s\geq 0$, under the assumption of small mass initial data in the focusing case [3], and further classified the traveling wave solutions of (1.1) and (1.3) on the circle [4]. Notice that a new numerical scheme based on explicit formulae was recently developped for (1.1) and (1.3) on the circle [2]. More studies on the intermediate nonlinear Schrödinger equation can be found in [26, 27, 28, 8].
In [21], Kim and Kwon established the soliton resolution results for both finite-time blow-up solutions and global-in-time solutions to (1.3). More precisely, they proved the scattering of solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R}):=\{u\in H^{1}(\mathbb{R}):xu\in L^{2}(\mathbb{R})\}$ with small mass initial data ($\|u_{0}\|_{L^{2}}^{2}<2\pi$) and the soliton resolution for global solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R})$. Following their approach, it is natural to expect that solutions to the defocusing (CM-DNLS) (1.1) in $H^{1,1}(\mathbb{R})$ also scatter as time tends to infinity.
In this paper, we applied the Gérard-type explicit formula to prove the scattering result for solutions to the defocusing (CM-DNLS) (1.1) with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. Our analysis also relies on the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$, and we can characterize the scattering term using this distorted Fourier transform. Following our approach developed in this paper, combined with the distorted Fourier transform corresponding to the focusing (CM-DNLS) (1.3) introduced in [11, Section 4], we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) (1.3) in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$.
1.1 The Lax pair
In this paper, we denote by $L_{+}^{2}(\mathbb{R})$ the Hardy space corresponding to $L^{2}(\mathbb{R})$ functions having a Fourier transform supported in the domain $\xi\geq 0$. Also, we denote by $L_{+}^{p}(\mathbb{R})$($1\leq p\leq\infty$) the space corresponding to $L^{p}(\mathbb{R})$ functions having a Fourier transform supported in the domain $\xi\geq 0$. Recall that the space $L_{+}^{2}(\mathbb{R})$ identifies to holomorphic functions on the upper-half plane $\mathbb{C}_{+}:=\{z\in\mathbb{C}:\operatorname{Im}(z)>0\}$ such that
$\sup_{y>0}\int_{\mathbb{R}}|f(x+iy)|^{2}dx<+\infty.$
The Riesz-Szegő projector $\Pi$ is the orthogonal projector from $L^{2}(\mathbb{R})$ onto $L_{+}^{2}(\mathbb{R})$. It is given by
$\forall f\in L^{2}(\mathbb{R}),\quad\forall z\in\mathbb{C}_{+},\quad\Pi f(z)=\frac{1}{2\pi}\int_{0}^{\infty}\mathrm{e}^{iz\xi}\widehat{f}(\xi)d\xi=\frac{1}{2i\pi}\int_{\mathbb{R}}\frac{f(y)}{y-z}dy.$ (1.4)
The Toeplitz operator $T_{b}$ on $L^{2}(\mathbb{R})$ associated to a function $b\in L^{\infty}(\mathbb{R})$ is defined by
$T_{b}f:=\Pi(bf),\qquad\forall f\in L_{+}^{2}(\mathbb{R}).$
For $u\in H_{+}^{1}(\mathbb{R})$, the Lax operator $L_{u}$ corresponding to (1.1) is defined by
$\forall f\in\operatorname{Dom}\left(L_{u}\right)=H_{+}^{1}(\mathbb{R}),\qquad L_{u}f:=Df+uT_{\bar{u}}f\text{ with }D:=\frac{1}{i}\frac{d}{dx}.$
For $u\in L_{+}^{2}(\mathbb{R})$, the Lax operator $L_{u}$ can be defined via an approach using quadratic forms with $\operatorname{Dom}(L_{u})\subset H_{+}^{\frac{1}{2}}(\mathbb{R})$ (see also [12, Appendix A]).
For $u\in H_{+}^{2}(\mathbb{R})$, another operator in the Lax pair, denoted by $B_{u}$, is defined by
$B_{u}=-uT_{\partial_{x}\bar{u}}+\partial_{x}uT_{\bar{u}}+i\left(uT_{\bar{u}}\right)^{2}.$
Also, we recall the definition of $X^{*}$ in [15],
$\forall f\in\operatorname{Dom}\left(X^{*}\right):=\left\{f\in L_{+}^{2}(\mathbb{R}):\exists\lambda_{f}\in\mathbb{C}:xf+\lambda_{f}\in L^{2}(\mathbb{R})\right\}=\left\{f\in L_{+}^{2}(\mathbb{R}):\widehat{f}\in H^{1}(0,\infty)\right\},$ (1.5)
with
$\widehat{X^{*}f}(\xi):=i\frac{d}{d\xi}[\widehat{f}(\xi)]\mathbf{1}_{\xi>0}.$
Here $X^{*}$ is the adjoint of the operator of multiplication by $x$ on $L_{+}^{2}(\mathbb{R})$, and $\left(\operatorname{Dom}\left(X^{*}\right),-iX^{*}\right)$ is maximally dissipative. Notice that $iX^{*}$ is the infinitesimal generator of the adjoint semi–group of contractions $\left(S(\eta)^{*}\right)_{\eta\geq 0}$ with
$\forall\eta\geq 0,\qquad S(\eta)^{*}=\mathrm{e}^{-i\eta X^{*}}.$
We also notice that
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\left|\widehat{f}\left(0_{+}\right)\right|^{2}=-4\pi\operatorname{Im}\left\langle X^{*}f,f\right\rangle_{L^{2}(\mathbb{R})}\leq 4\pi\|X^{*}f\|_{L^{2}}\|f\|_{L^{2}}.$
Therefore, we can define
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\quad I_{+}(f):=\widehat{f}\left(0_{+}\right).$ (1.6)
Moreover, we can easily deduce that
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\quad X^{*}f=xf+\frac{1}{2i\pi}I_{+}(f).$ (1.7)
1.2 Maximally dissipative operators and the Kato-Rellich theorem
In this paper, we will also apply the Miyadera–Gustafson theorem to show that operators remain maximally dissipative after some perturbations. To present the Miyadera–Gustafson theorem for maximally dissipative operators, we first need to introduce the following definition of the dissipative and maximally dissipative operators in Banach spaces.
Definition 1.1.
Let $(D(A),A)$ be an operator in a Banach space $\mathscr{B}$.
1. We say that $A$ is dissipative if for all $g\in D(A)$ and all $\lambda>0$,
$\|(\lambda I-A)g\|\geq\lambda\|g\|.$
2. We say that $A$ is maximal dissipative if it is dissipative and for all $h\in\mathscr{B}$ and for all $\lambda>0$, there exists $g\in D(A)$ such that $(\lambda I-A)g=h$.
Remark 1.2.
In fact, an operator $(D(A),A)$ in a Hilbert space $\mathscr{H}$ is dissipative if and only if for all $g\in D(A),\,\Re\langle Ag,g\rangle\leq 0$.
Remark 1.3.
Let $(D(A),A)$ be a maximally dissipative operator in a Banach space $\mathscr{B}$. From the definition of maximally dissipative operators, we can deduce that, for all $\lambda>0$, we have
$\|(\lambda I-A)^{-1}\|_{\mathscr{L}(\mathscr{B})}\leq\frac{1}{\lambda},\qquad\|A(\lambda I-A)^{-1}\|_{\mathscr{L}(\mathscr{B})}\leq 2.$ (1.8)
Since the Miyadera–Gustafson theorem involves the related notion of the perturbation of operators, we give the following definition of the relative bound of an operator with respect to another operator.
Definition 1.4.
Let $\left(D(A),A\right)$ and $\left(D(B),B\right)$ be densely defined linear operators on a Banach space $\mathscr{B}$. Suppose that:
(i) $D(A)\subset D(B)$;
(ii) For some $a$ and $b$ in $\mathbb{R}$ and all $\varphi\in D(A)$,
$\|B\varphi\|\leq a\|A\varphi\|+b\|\varphi\|.$
Then $B$ is said to be $A$-bounded. The infimum of such $a$ is called the relative bound of $B$ with respect to $A$. If the relative bound is 0, we say that $B$ is infinitesimally small with respect to $A$.
Then we state the Miyadera–Gustafson theorem for maximally dissipative operators.
Theorem 1.5 (Miyadera–Gustafson theorem).
Let $\left(D(A),A\right)$ be a maximally dissipative operator which is densely defined on a Banach space $\mathscr{B}$ and assume $\left(D(B),B\right)$ to be dissipative and $A$-bounded with the relative bound smaller than 1. Then $\left(D(A),A+B\right)$ is also a maximally dissipative operator.
We refer to [25] for the proof of the Miyadera–Gustafson theorem.
1.3 Global well-posedness
We have the following global-well-posedness result for the defocusing (CM-DNLS) (1.1).
Theorem 1.6 ([12, 19]).
Given $u_{0}\in H_{+}^{s}(\mathbb{R})$ with $s\geq 0$, there exists a solution $u\in C(\mathbb{R};H_{+}^{s}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0,x)=u_{0}(x)$.
Remark 1.7.
We have also the analogous global well-posedness result in $H_{+}^{s}(\mathbb{R}))$ ($s\geq 0$) for the focusing (CM-DNLS) (1.3) with small mass initial data ($\|u\|_{L^{2}}^{2}<2\pi$, or $u_{0}\in H_{+}^{1}(\mathbb{R})$ with $\|u_{0}\|_{L^{2}}^{2}=2\pi$) [12, 19]. For the focusing (CM-DNLS) (1.3) with large mass initial data ($\|u\|_{L^{2}}^{2}>2\pi$), the finite-time blow up phenomena would occur [20].
1.4 Gérard-type explicit formula
In [19], Killip, Laurens and Vişan followed Gérard’s approach used in [15] to derive an explicit formula for the solution to (1.1) in $L_{+}^{2}(\mathbb{R})$.
Theorem 1.8.
([19, Theorem 1.7]) Let $u\in C\left(\mathbb{R},L_{+}^{2}(\mathbb{R})\right)$ be the solution to (1.1) with $u(0)=u_{0}$. Then for any $t\in\mathbb{R}$,
$\forall z\in\mathbb{C}_{+}:=\{z\in\mathbb{C}:\,\Im(z)>0\},\quad u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right].$ (1.9)
The Gérard-type explicit formula was first founded by Gérard in [15] for the Benjamin–Ono equation on the line (see also [9] for the extension), and then by several mathematicians for the cubic Szegő equation on the line [14], the Calogero–Moser DNLS equation [19, 10], and the half wave maps equation [13]. The Gérard-type explicit formula has also been used in several studies [14, 16, 19, 9, 5, 6, 7, 13] so far and has shown to be very effective.
1.5 Main result
In this paper, we used the Gérard-type explicit formula (1.9) to study the long-time behavior of (1.1). More precisely, we proved the following scattering result for solutions to (1.1) with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u_{0}\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u_{0}\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$.
Theorem 1.9.
Let $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R}):=\{w\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. Then the solution $u\in C(\mathbb{R};L_{+}^{2}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
$\lim_{t\rightarrow+\infty}\left\|u(t)-\mathrm{e}^{i\partial_{x}^{2}}v_{0}\right\|_{L^{2}(\mathbb{R})}=0.$ (1.10)
Here $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$, and
$\mathcal{F}_{d}(f)=\widetilde{f},\qquad\forall f\in L_{+}^{2}(\mathbb{R})$
is the distorted Fourier transform assciated with $L_{u_{0}}$ (see Section 2.1 and Section 2.2 for details).
Remark 1.10.
As mentioned in the introduction, Kim and Kwon proved the scattering result for solutions to (1.3) in $H^{1,1}(\mathbb{R})$ [21]. More precisely, their appraoch relies on the pseudo-conformal symmetry, and thus requires $xu_{0}\in L^{2}(\mathbb{R})$. Also, it is expected that we can follow their approach to show the scattering result for the defocusing (CM-DNLS) (1.1) in $H^{1,1}(\mathbb{R})$. The great progress we made in this paper is to extend the chiral scattering result to more singular chiral initial data in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$. On the other hand, we can characterize the scattering term using the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$ of the defocusing (CM-DNLS) (1.1). In fact, following the method used in this paper, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [11, Section 4], we can also establish the scattering result for solutions to the focusing (CM-DNLS) (1.3) with initial data in $L_{+}^{2,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ with some $\alpha>0$, which is an extension of the chiral scattering result for the focusing (CM-DNLS) (1.3) in [21]. See Remark 3.2 for details. Another highlight of this paper is that it is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. See also [6] for the study on the long-time behavior of the Benjamin–Ono equation with rational initial data and [13] for the study on the long-time behavior of the half-wave maps equation with rational initial data.
Remark 1.11.
In [21], Kim and Kwon also showed the soliton resolution for global solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R})$. With a slight modification of the proof of Theorem 1.9, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [11, Section 4], we can conclude the asymptotic bound-state/radiation decomposition for global solutions to (1.3) in $L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$, i.e., if $u\in C(\mathbb{R};L_{+}^{2}(\mathbb{R}))$ is a global solution to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ for some $\alpha>0$, then
$u(t,x)=\text{bound-states}+\mathrm{e}^{i\partial_{x}^{2}}v_{0}+o(1)\text{ as }t\to\infty$
in $L_{+}^{2}(\mathbb{R})$. Here $v_{0}$ can be defined as in Theorem 1.9, using the distorted Fourier transform introduced in [11, Section 4]. To establish soliton resolution in this setting, one would need to show that each bound state asymptotically converges to a traveling soliton profile. See Remark 3.2 and Remark 3.3 for details.
Remark 1.12.
One can observe that the (unique) ground state solution $\frac{\sqrt{2}}{x+i}$ to the focusing (CM-DNLS) (1.3) (see [12, Lemma 4.1]) is in $L_{+}^{2,\alpha}(\mathbb{R})$ for $0<\alpha<\frac{1}{2}$ but not in $H^{1,1}(\mathbb{R})$. This shows that our progress on the long-time analysis of (1.1) and (1.3) is crucial.
2 Preliminaries
In this section, we study the generalized eigenfunctions and the distorted Fourier transform associated with the Lax operator $L_{u}$ of (1.1). For $u\in L_{+}^{2}(\mathbb{R})$, it is known that $L_{u}$ has no eigenvalue in this case. Under some additional decay assumption on $u\in L_{+}^{2}(\mathbb{R})$, we show the existence of the generalized eigenfunctions in $L_{+}^{\infty}(\mathbb{R})$ and then introduce the corresponding distorted Fourier transform. See also [11, Section 4] for the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3).
2.1 Generalized eigenfunctions
First we show that if $u\in L_{+}^{2}(\mathbb{R})$, the Lax operator $L_{u}$ has no eigenvalue.
Proposition 2.1.
Let $u\in L_{+}^{2}(\mathbb{R})$, then $L_{u}$ has no eigenvalues.
Proof.
Let $\psi\in H_{+}^{\frac{1}{2}}(\mathbb{R})$ be an eigenvector of $L_{u}$. From the proof of [12, Proposition 5.1], we can infer
$\|\psi\|_{L^{2}}^{2}+\frac{1}{2\pi}|\langle\psi,u\rangle|^{2}=0.$
Thus $\psi=0$. ∎
Remark 2.2.
Using the Weyl’s theorem, we can easily conclude that $\sigma_{ess}(L_{u})=\sigma_{ess}(D)=[0,\infty)$. Since $L_{u}$ has no eigenvalues, we can conclude that $\sigma(L_{u})=\sigma_{ac}(L_{u})\cup\sigma_{sc}(L_{u})=[0,\infty)$.
Remark 2.3.
For the Lax operator of the focusing (CM-DNLS) (1.3), if $\|u\|_{L^{2}}^{2}<2\pi$, we can also infer that this Lax operator has no eigenvalues.
We now give a fundamental result concerning the generalized eigenfunctions of $L_{u}$, which constitute the foundation of the distorted Fourier transform.
Proposition 2.4.
Let $u\in L^{2,\alpha}_{+}(\mathbb{R})$ with some $\alpha>0$. For every $\lambda>0$, the space
$\mathscr{M}_{u}(\lambda)=\bigl\{\,m\in L^{\infty}_{+}(\mathbb{R}):\;(L_{u}-\lambda)m=0\ \text{in the distributional sense}\bigr\}$
is one–dimensional. Every element $m\in\mathcal{M}_{u}(\lambda)$ satisfies
$\mathrm{e}^{-i\lambda x}m(x).$
has a limit $\ell_{m}(\pm\infty)\text{ as }x\rightarrow\pm\infty$. Furthermore, $\ell_{m}(-\infty)$ characterizes $m$ and
$\left|\ell_{m}(+\infty)\right|=\left|\ell_{m}(-\infty)\right|.$
Proof.
Notice that, if $m\in L^{\infty}(\mathbb{R})$,
$\Pi(\bar{u}m)\in L_{+}^{2}(\mathbb{R}),$
so that $uT_{\bar{u}}m\in L^{1}(\mathbb{R})$. Equation $(L_{u}-\lambda)m=0$ is equivalent to
$D\left(\mathrm{e}^{-i\lambda x}m(x)\right)=-\mathrm{e}^{-i\lambda x}u(x)\Pi(\bar{u}m)(x),$
which implies that
$D\left(\mathrm{e}^{-i\lambda x}m(x)\right)\in L^{1}(\mathbb{R}),$
hence $\mathrm{e}^{-i\lambda x}m(x)$ has a limit $\ell_{m}(\pm\infty)$ as $x\rightarrow\pm\infty$. Furthermore,
$i\left(\left|\ell_{m}(+\infty)\right|^{2}-\left|\ell_{m}(-\infty)\right|^{2}\right)$ $=\lim_{R\rightarrow+\infty}\int_{-R}^{R}[m(x)\overline{Dm(x)}-Dm(x)\overline{m(x)}]dx$
$=\int_{-\infty}^{\infty}\left[-m(x)\overline{uT_{\bar{u}}m(x)}+uT_{\bar{u}}m(x)\overline{m(x)}\right]dx$
$=-\left\|T_{\bar{u}}m\right\|^{2}+\left\|T_{\bar{u}}m\right\|^{2}$
$=0.$
This proves $\left|\ell_{m}(+\infty)\right|=\left|\ell_{m}(-\infty)\right|$. It remains to prove that the space of such functions $m$ is one dimensional.
Given $\ell\in\mathbb{C}$, the problem $\left(L_{u}-\lambda\right)m=0,\ell_{m}(-\infty)=\ell$ is equivalent to
$m=\mathrm{e}^{i\lambda x}a,\quad a-K_{\lambda}(a)=\ell,$ (2.1)
where $K_{\lambda}:L^{\infty}(\mathbb{R})\rightarrow L^{\infty}(\mathbb{R})$ is defined as
$K_{\lambda}(a)(x):=-i\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}a\right)(y)dy.$
We claim that $1-K_{\lambda}:L^{\infty}\rightarrow L^{\infty}$ is bijective. We shall use the Fredholm alternative by showing that $K_{\lambda}$ is compact and $1-K_{\lambda}$ is injective.
Lemma 2.5.
The operator $K_{\lambda}:L^{\infty}(\mathbb{R})\rightarrow L^{\infty}(\mathbb{R})$ is compact.
Proof.
If $\|a\|_{L^{\infty}}\leq 1$, $\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}a\right)$ belongs to a bounded subset of $L^{1}$. Furthermore,
$\left|K_{\lambda}(a)(x)-K_{\lambda}(a)\left(x^{\prime}\right)\right|\leq\|u\|_{L^{2}}\left(\int_{x}^{x^{\prime}}|u(y)|^{2}dy\right)^{\frac{1}{2}},\qquad x\leq x^{\prime}$
$\left|K_{\lambda}(a)(x)\right|\leq\|u\|_{L^{2}}\left(\int_{-\infty}^{x}|u(x)|^{2}dx\right)^{\frac{1}{2}}$
$\left|K_{\lambda}(a)(x)-K_{\lambda}(a)(+\infty)\right|\leq\|u\|_{L^{2}}\left(\int_{x}^{+\infty}|u(x)|^{2}dx\right)^{\frac{1}{2}}$
For all these reasons, the set
$\left\{K_{\lambda}(a),\|a\|_{L^{\infty}}\leq 1\right\}$
is relatively compact in $L^{\infty}$. ∎
Lemma 2.6.
The operator $\operatorname{Id}-K_{\lambda}:L^{\infty}\rightarrow L^{\infty}$ is injective.
Proof.
Let $f\in L^{\infty}$ such that $f=K_{\lambda}(f)$. Let us prove that $f\in L^{2}(\mathbb{R})$. First, we observe that
$m(x):=\mathrm{e}^{i\lambda x}f(x)$
satisfies $\left(L_{u}-\lambda\right)m=0$ and $\ell_{m}(-\infty)=0$, therefore $\ell_{m}(+\infty)=0$, or equivalently
$\int_{-\infty}^{+\infty}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)dy=0,$
so that
$f(x)=\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)dy=-\int_{x}^{+\infty}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)dy.$
Let $\alpha>0$ such that $|x|^{\alpha}u\in L^{2}$. Then
$g(x):=\bar{u}(x)\mathrm{e}^{i\lambda x}f(x)$
satisfies $\left(1+|x|^{\alpha}\right)g\in L^{2}(\mathbb{R})$, which is equivalent to $\widehat{g}\in H^{\alpha}(\mathbb{R})$. By Hardy’s inequality, $\mathbf{1}_{\mathbb{R}_{+}}\widehat{g}\in H^{\alpha}(\mathbb{R})$, or $\left(1+|x|^{\alpha}\right)\Pi g\in L^{2}(\mathbb{R})$. This implies
$|y|^{2\alpha}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda y}f\right)(y)\in L^{1}(\mathbb{R}),$
and consequently,
$|x|^{2\alpha}|f(x)|\in L^{\infty}(\mathbb{R}).$
If $\alpha>\frac{1}{4}$, this implies $f\in L^{2}(\mathbb{R})$. Otherwise, we plug this information in the above argument. More generally, if $|x|^{\beta}f\in L^{\infty}$ for some $\beta\geq 0$, we have $\left(1+|x|^{\alpha+\beta}\right)f\in L^{2}(\mathbb{R})$, hence, again by Hardy’s inequality,
$\left(1+|x|^{\gamma}\right)g\in L^{2}(\mathbb{R}),$
with $\gamma=\alpha+\beta$ if $\alpha+\beta<\frac{1}{2}$, and $\gamma=\frac{1}{2}-\varepsilon$ for every $\varepsilon>0$ if $\alpha+\beta\geq\frac{1}{2}$. Consequently,
$|x|^{\alpha+\gamma}f\in L^{\infty}$
which leads to $f\in L^{2}$ after a finite number of iterations. Since
$m(x)=\mathrm{e}^{i\lambda x}f(x),$
we infer that $m\in L_{+}^{2}\text{, and finally }m\in H_{+}^{\frac{1}{2}}$ is an eigenvector of $L_{u_{0}}$. Since the kernel of $L_{u}-\lambda$ is trivial, we conclude $m=0$. ∎
Now we can conclude the proof of Proposition 2.4. By the Fredholm alternative, the operator $\operatorname{Id}-K_{\lambda}:L^{\infty}(\mathbb{R})\rightarrow L^{\infty}(\mathbb{R})$ is bejective. Coming back to (2.1), the linear form
$m\in\mathscr{M}_{u}(\lambda)\mapsto\ell_{m}(-\infty)\in\mathbb{C}$
is bijective. Therefore $\mathscr{M}_{u}(\lambda)$ is one dimensional. ∎
If $u\in\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$, we can also conclude the limit of $m(x,\lambda)$ as $\lambda\to 0_{+}$, which is useful in the latter proof.
Lemma 2.7.
Let $u\in\mathcal{S}_{+}(\mathbb{R})$ and $\ell_{m}(-\infty)=1$. For any $x\in\mathbb{R}$, we have
$\lim_{\lambda\to 0_{+}}m(x,\lambda)=1.$
Proof.
Let $\lambda=0$ in (2.1), then following the previous operations, we can also prove that the solution to (2.1) with $\lambda=0$ exists and is unique. It is not hard to see that $m=a=1$ is the solution to (2.1) with $\lambda=0$ and $\ell=1$. Now we define $K_{0}:=K_{\lambda}|_{\lambda=0}$, and we can easily infer that
$\forall f\in L^{\infty}(\mathbb{R}),\qquad\lim_{\lambda\to 0_{+}}\|(K_{\lambda}-K_{0})f\|_{L^{\infty}}=0.$
Then from the second resolvent identity, we can deduce that
$\left\|a(\cdot,\lambda)-1\right\|_{L^{\infty}}\leq\left\|(\operatorname{Id}-K_{\lambda})^{-1}1-(\operatorname{Id}-K_{0})^{-1}1)\right\|_{L^{\infty}}\rightarrow 0\text{ as }\lambda\to 0_{+}.$
Since for any $x\in\mathbb{R}$, $\mathrm{e}^{i\lambda x}\to 1$ as $\lambda\to 0_{+}$, we can conclude the lemma. ∎
2.2 The distorted Fourier transform
From Proposition 2.4, for $u\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$, there exists a unique $m_{-}\in\mathscr{M}_{u}(\lambda)$ such that $\ell_{m_{-}}(-\infty)=1$. Given $f\in L_{+}^{1}(\mathbb{R})$, we can define
$\mathcal{F}_{d}f(\lambda):=\widetilde{f}(\lambda):=\int_{\mathbb{R}}f(x)\overline{m_{-}(x,\lambda)}dx,\qquad\lambda>0.$
If moreover $f\in L^{2}(\mathbb{R})$, then $\widetilde{f}\in L^{2}((0,+\infty))$. Denote by $\mathcal{H}_{c}\left(L_{u}\right)$ the continuous space of $L_{u}$, namely the space of those $f\in L_{+}^{2}(\mathbb{R})$ having a spectral measure with no pure point part. Denote by $P_{c}(L_{u})$ the orthogonal projector from $L_{+}^{2}$ to $\mathcal{H}_{c}\left(L_{u}\right)$.
Proposition 2.8.
The operator $f\mapsto\widetilde{f}$ extends to $L_{+}^{2}(\mathbb{R})$ with the identity
$\int_{\mathbb{R}}\left|f(x)\right|^{2}dx=\int_{\mathbb{R}}\left|P_{c}(L_{u})f(x)\right|^{2}dx=\frac{1}{2\pi}\int_{0}^{\infty}|\widetilde{f}(\lambda)|^{2}d\lambda.$ (2.2)
Proof.
From the spectral theorem and the fact that $L_{u}$ has no eigenvalue, we have, for every $f\in L_{+}^{2}(\mathbb{R})$,
$\|f\|_{L^{2}}^{2}=\left\|P_{c}(L_{u})f\right\|^{2}=\frac{1}{2i\pi}\lim_{\varepsilon\rightarrow 0_{+}}\int_{0}^{\infty}\left\langle\left(L_{u}-\lambda-i\varepsilon\right)^{-1}f-\left(L_{u}-\lambda+i\varepsilon\right)^{-1}f,f\right\rangle d\lambda.$ (2.3)
Set $g_{\pm}^{\varepsilon}(.,\lambda):=\left(L_{u}-\lambda\mp i\varepsilon\right)^{-1}f$. We have
$g_{\pm}^{\varepsilon}=(D-\lambda\mp i\varepsilon)^{-1}\left(Id+uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}\right)^{-1}f.$
The following lemma can be proved along the same lines as the properties of operator $K_{\lambda}$ in the previous section.
Lemma 2.9.
The following properties hold for every $\lambda\geq 0$, uniformly if $\lambda$ belongs to a compact set of $0,\infty)$.
1. $(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{\infty}$ is uniformly bounded and strongly convergent to $(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{\infty}$ as $\varepsilon\rightarrow 0_{+}$.
2. $uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{1}$ is uniformly compact as $\varepsilon\rightarrow 0_{+}$, and is convergent to $uT_{\bar{u}}(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{1}$ as $\varepsilon\rightarrow 0_{+}$.
3. $\left(Id+uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}\right)^{-1}:L^{1}\rightarrow L^{1}$ is strongly convergent to
$\left(Id+uT_{\bar{u}}(D-\lambda\mp i0)^{-1}\right)^{-1}:L^{1}\rightarrow L^{1}$
as $\varepsilon\rightarrow 0_{+}$.
Assume that $f\in L_{+}^{1}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})$. From this lemma, we infer that $g_{\pm}^{\varepsilon}$ tends to $g_{\pm}$ in $L^{\infty}$ with
$g_{+}(x,\lambda)=\int_{-\infty}^{x}\mathrm{e}^{i\lambda(x-y)}f(y)dy-\int_{-\infty}^{x}\mathrm{e}^{i\lambda(x-y)}u(y)\Pi_{+}\left(\bar{u}g_{+}(.,\lambda)\right)(y)dy$
$g_{-}(x,\lambda)=-\int_{x}^{+\infty}\mathrm{e}^{i\lambda(x-y)}f(y)dy+\int_{x}^{+\infty}\mathrm{e}^{i\lambda(x-y)}u(y)\Pi_{+}\left(\bar{u}g_{-}(.,\lambda)\right)(y)dy$
In particular, from Proposition 2.4, $g_{+}-g_{-}=-\ell_{g_{-}}(-\infty)m_{-}$, with
$-i\ell_{g_{-}}(-\infty)=\lim_{R\rightarrow+\infty}\int_{-R}^{R}\left(g_{-}\overline{Dm_{-}}-Dg_{-}\bar{m}_{-}\right)dx=-\int_{-\infty}^{\infty}f\bar{m}_{-}dx=-\widetilde{f}(\lambda)$
and finally
$g_{+}(x,\lambda)-g_{-}(x,\lambda)=i\widetilde{f}(\lambda)m_{-}(x,\lambda)$
and, passing to the limit in the right hand side of (2.3) leads to (2.2) for $f\in L_{+}^{1}\cap L_{+}^{2}$. By density, the operator $f\mapsto\widetilde{f}$ extends to $L_{+}^{2}(\mathbb{R})$, with identity (2.2). ∎
Remark 2.10.
By the polarization identity, the isometry property implies preservation of the inner product, i.e., for any $f,g\in L_{+}^{2}(\mathbb{R})$, we have
$\left\langle f,g\right\rangle_{L^{2}(\mathbb{R})}=\frac{1}{2\pi}\langle\widetilde{f},\widetilde{g}\rangle_{L^{2}(0,\infty)}.$ (2.4)
Remark 2.11.
Notice that, if $f\in H_{+}^{1}(\mathbb{R})$,
$\widetilde{L_{u}f}(\lambda)=\int_{\mathbb{R}}L_{u}f(x)\overline{m_{-}(x,\lambda)}dx=\int_{\mathbb{R}}f(x)\overline{L_{u}m_{-}(x,\lambda)}dx=\lambda\widetilde{f}(\lambda).$
Applying the spectral theorem to the self-adjoint operator $L_{u}$ restricted to $\mathcal{H}_{c}(L_{u})$, combined with Proposition 2.4, we can infer the surjectivity of the distorted Fourier transform.
Corollary 2.12.
The map $f\in L_{+}^{2}(\mathbb{R})\mapsto\widetilde{f}\in L^{2}(0,\infty)$ is onto.
Remark 2.13.
From the the spectral theorem and Proposition 2.8, we can also conclude that $\sigma(L_{u})=\sigma_{ac}(L_{u})=[0,\infty)$.
Using the interpolation between $\mathcal{F}_{d}:L_{+}^{1}(\mathbb{R})\to L^{\infty}(0,\infty)$ and $\mathcal{F}_{d}:L_{+}^{2}(\mathbb{R})\to L^{2}(0,\infty)$, we can also extend the distorted Fourier transform $\mathcal{F}_{d}$ to $\mathcal{F}_{d}:L_{+}^{p}\left(\mathbb{R}\right)\rightarrow L^{p^{\prime}}\left(0,\infty\right)$ for $1<p<2$.
Corollary 2.14.
The distorted Fourier transform can be extended uniquely and continuously to
$\mathcal{F}_{d}:L_{+}^{p}\left(\mathbb{R}\right)\rightarrow L^{p^{\prime}}\left(0,\infty\right),\qquad 1\leq p\leq 2,$
where
$\frac{1}{p}+\frac{1}{p^{\prime}}=1.$
As is well known, the classical Fourier transform is a continuous bijection from $\mathcal{S}(\mathbb{R})$ onto itself. Similarly, for the distorted Fourier transform defined above, if we restrict to $u\in\mathcal{S}_{+}(\mathbb{R})$, we obtain an analogous property.
Proposition 2.15.
For $u\in\mathcal{S}_{+}(\mathbb{R})$, the distorted Fourier transform $\mathcal{F}_{d}:\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})\rightarrow\mathcal{S}_{*}:=\{f=\left.g\right|_{(0,\infty)}:g\in\mathcal{S}(\mathbb{R})\text{ and }\operatorname{supp}(g)\subset[0,\infty)\}$ is well-defined and a continuous bijection.
Proof.
Let us recall (2.1) with $\ell=1$,
$m_{-}=\mathrm{e}^{i\lambda x}a,\qquad a-K_{\lambda}(a)=1\text{ with }K_{\lambda}(a)(x):=-i\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}a\right)(y)dy.$
so we have
$a=(Id-K_{\lambda})^{-1}(1).$
Then
$\partial_{x}a(x,\lambda)=-i\mathrm{e}^{-i\lambda x}u(x)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}a\right)(x)\in L_{x,\lambda}^{\infty}.$
From the above expression, it is not hard to see that
$\forall n\in\mathbb{N}_{+},\qquad\partial_{x}^{n}a(x,\lambda)\in L_{x,\lambda}^{\infty}.$
Also,
$\partial_{\lambda}a(x,\lambda)$ $=(Id-K_{\lambda})^{-1}\partial_{\lambda}K_{\lambda}(Id-K_{\lambda})^{-1}(1)$
$=(Id-K_{\lambda})^{-1}\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u[T_{\bar{u}},y]\mathrm{e}^{i\lambda y}a(y,\lambda)dy$
$=\frac{1}{2i\pi}(Id-K_{\lambda})^{-1}\left(\bar{\widetilde{u}}(\lambda)\int_{-\infty}^{x}u(y)\mathrm{e}^{-i\lambda y}dy\right),$
here the last identity comes from
$\forall f\in L^{\infty},\qquad u[T_{\bar{u}},y]f=\frac{u}{2i\pi}\int_{\mathbb{R}}f\bar{u}dx.$
Since
$\left|\lambda\widetilde{u}(\lambda)\right|$ $=\left|\lambda\int_{\mathbb{R}}u(x)\mathrm{e}^{-ix\lambda}\bar{a}(x,\lambda)dx\right|$
$=\left|\int_{\mathbb{R}}u(x)\partial_{x}(\mathrm{e}^{-ix\lambda})\bar{a}(x,\lambda)dx\right|$
$\leq\left|\int_{\mathbb{R}}u^{\prime}(x)\mathrm{e}^{-ix\lambda}\bar{a}(x,\lambda)dx\right|+\left|\int_{\mathbb{R}}u(x)\mathrm{e}^{-ix\lambda}\partial_{x}\bar{a}(x,\lambda)dx\right|$
$\leq C,$
we deduce that
$\|\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-1}.$
Furthermore, it is not hard to conclude
$\forall k\in\mathbb{N}_{+},\qquad\|\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq C_{k}(1+|\lambda|)^{-k}.$
Moreover, since
$\partial_{x}\partial_{\lambda}a=\frac{1}{2i\pi}\widetilde{u}(\lambda)u(x)\mathrm{e}^{-i\lambda x}-i\mathrm{e}^{-i\lambda x}u(x)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}\partial_{\lambda}a\right)(x),$
we have
$\|\partial_{x}\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq\frac{1}{2\pi}\|\widetilde{u}(\lambda)u(x)\|_{L_{x}^{\infty}}+\|\mathrm{e}^{-i\lambda x}u(x)\Pi\left(\bar{u}\mathrm{e}^{i\lambda\cdot}\partial_{\lambda}a\right)(x)\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-1}.$
It is not hard to conclude
$\forall n\in\mathbb{N}_{+},\qquad\|\partial_{x}^{n}\partial_{\lambda}a\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-1}.$
For higher $\lambda$–derivatives, differentiating repeatedly in $\lambda$ produces sums of terms where derivatives fall on $\mathrm{e}^{-i\lambda y},\mathrm{e}^{i\lambda y}$ and on $a$. Each time we gain a primitive like $\int_{-\infty}^{x}y^{k}u(y)e^{-i\lambda y}dy$ (which decays faster than any power of $|\lambda|$ since $u\in\mathcal{S}_{+}(\mathbb{R})$), $\int_{\mathbb{R}}y^{k}\bar{u}(y)e^{i\lambda y}\partial_{\lambda}^{\ell}a(y,\lambda)dy$ or $\partial_{\lambda}^{k}\bar{\widetilde{u}}(\lambda)$. Then we can follow the previous argument and use the recurrence to show that
$\forall n\in\mathbb{N}_{+}\text{ and }\forall m\in\mathbb{N}_{+},\qquad\|\partial_{x}^{n}\partial_{\lambda}^{m}a\|_{L_{x}^{\infty}}\leq C(1+|\lambda|)^{-m}.$
So $a$ is a symbol of order 0, by the classical pseudo-differential operators theory, we can conclude that the distorted Fourier transform $\mathcal{F}_{d}:\mathcal{S}_{+}(\mathbb{R}):=\mathcal{S}(\mathbb{R})\cap L_{+}^{2}(\mathbb{R})\rightarrow\mathcal{S}_{*}:=\{f=\left.g\right|_{(0,\infty)}:g\in\mathcal{S}(\mathbb{R})\text{ and }\operatorname{supp}(g)\subset[0,\infty)\}$ is well-defined and a continuous map. Proposition 2.8 implies the injection of $\mathcal{F}_{d}$. Since the distorted Fourier transform is a unitary operator, we can easily infer the surjection of $\mathcal{F}_{d}$ from $\mathcal{S}_{+}(\mathbb{R})$ to $\mathcal{S}_{*}$. ∎
Corollary 2.16.
For $u\in\mathcal{S}_{+}(\mathbb{R})$, the distorted Fourier transform $\mathcal{F}_{d}$ extends uniquely and continuously to the tempered distributions: $\mathcal{F}_{d}:\mathcal{S}_{+}^{\prime}(\mathbb{R})\to\mathcal{S}_{*}^{\prime}(\mathbb{R})$, via
$\langle\mathcal{F}_{d}F,\ \phi\rangle:=2\pi\langle F,\ \mathcal{F}_{d}^{-1}\phi\rangle,\qquad\forall F\in\mathcal{S}^{\prime}_{+}(\mathbb{R}),\phi\in\mathcal{S}_{*}(\mathbb{R}).$
Proof.
By Proposition 2.15 we know that the operator
$\mathcal{F}_{d}f(\lambda)=\int_{\mathbb{R}}e^{-i\lambda x}\overline{a(x,\lambda)}f(x)\,dx$
maps $\mathcal{S}_{+}(\mathbb{R})$ continuously onto $\mathcal{S}_{*}$. Its inverse on $\mathcal{S}_{*}$ is given by
$\mathcal{F}_{d}^{-1}\varphi(x)=\frac{1}{2\pi}\int_{0}^{\infty}e^{i\lambda x}a(x,\lambda)\,\varphi(\lambda)\,d\lambda,$
which is continuous $\mathcal{S}_{*}\to\mathcal{S}_{+}$ and satisfies $\mathcal{F}_{d}^{-1}\circ\mathcal{F}_{d}=\mathrm{Id}_{\mathcal{S}_{+}}$ and $\mathcal{F}_{d}\circ\mathcal{F}_{d}^{-1}=\mathrm{Id}_{\mathcal{S}_{*}}$.
For $F\in\mathcal{S}_{+}^{\prime}(\mathbb{R})$, we can then define the distorted Fourier transform of this distribution by
$\langle\mathcal{F}_{d}F,\ \phi\rangle:=2\pi\langle F,\ \mathcal{F}_{d}^{-1}\phi\rangle,\qquad\forall F\in\mathcal{S}^{\prime}_{+}(\mathbb{R}),\phi\in\mathcal{S}_{*}(\mathbb{R}).$
Here $\mathcal{F}_{d}F\in\mathcal{S}_{*}^{\prime}(\mathbb{R})$. ∎
2.3 Computation of the scattering term
We have already completed the distorted Fourier transform theory in the previous part. In this section, using the distorted Fourier transform, we try to find the relation between $X^{*}f$ and $i\partial_{\lambda}\widetilde{f}$, and then we can compute the scattering term. First we go to analyze $\partial_{\lambda}a$. We know that, the problem $(L_{u}-\lambda)m_{-}=0,\ell_{m_{-}}(-\infty)=1$ is equivalent to
$m(\lambda,x)=e^{i\lambda x}a(x,\lambda),\qquad a(x,\lambda)=1+K_{\lambda}(a)(x),$
where
$K_{\lambda}(a)(x):=-i\int_{-\infty}^{x}\mathrm{e}^{-i\lambda y}u(y)\Pi\left(\bar{u}(y)\mathrm{e}^{i\lambda y}a(\lambda,y)\right)dy.$
Again, we assume that $u\in\mathcal{S}_{+}(\mathbb{R})$. We have $\partial_{\lambda}a$ is well defined in $L_{+}^{\infty}$ and
$\partial_{\lambda}a=\frac{1}{2i\pi}\bar{\widetilde{u}}(\lambda)\int_{-\infty}^{x}e^{-i\lambda y}u(y)dy+K_{\lambda}(\partial_{\lambda}a).$
Denote $n(x,\lambda):=e^{i\lambda x}\partial_{\lambda}a$, we infer
$L_{u}n-\lambda n=-\frac{1}{2\pi}\bar{\widetilde{u}}(\lambda)u(x).$
Applying the distorted Fourier transform and Proposition 2.4, we can deduce that
$\widetilde{n}(\mu,\lambda)=-\frac{1}{2\pi}p.v.\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)}{\mu-\lambda}+c\delta(\mu-\lambda),$
i.e.,
$n(x,\lambda)=n_{1}(x,\lambda)+c\,m_{-}(x,\lambda),$ (2.5)
where
$\widetilde{n}_{1}(\mu,\lambda):=-\frac{1}{2\pi}p.v.\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)}{\mu-\lambda}$
and $c$ to be determined.
To obtain $c$, we need to give the aysmptotic behavior of $\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$ as $x\to$-$\infty$. For this purpose, we establish the following lemma.
Lemma 2.17.
Let $u\in\mathcal{S}_{+}(\mathbb{R})$, then we have
$\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$
Proof.
By the classical Fourier transform theory, we know that
$p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)e^{ix\mu}}{\mu-\lambda}d\mu$
$=p.v.\int_{0}^{\infty}\frac{|\widetilde{u}(\lambda)|^{2}e^{ix\mu}}{\mu-\lambda}d\mu+p.v.\int_{0}^{\infty}\bar{\widetilde{u}}(\lambda)\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}e^{ix\mu}d\mu$
$=e^{i\lambda x}|\widetilde{u}(\lambda)|^{2}\left[-\mathrm{Ci}(\lambda|x|)+i\operatorname{sgn}(x)\left(\frac{\pi}{2}+\mathrm{Si}(\lambda|x|)\right)\right]+p.v.\int_{0}^{\infty}\bar{\widetilde{u}}(\lambda)\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}e^{ix\mu}d\mu,$
where
$\operatorname{Si}(w)=\int_{0}^{w}\frac{\sin t}{t}dt,\qquad\operatorname{Ci}(w)=-\int_{w}^{\infty}\frac{\cos t}{t}dt.$
Since $\widetilde{u}\in\mathcal{S}_{*}$, we know that
$\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}\in L^{1}([0,\infty)),$
so we infer that
$p.v.\int_{0}^{\infty}\bar{\widetilde{u}}(\lambda)\frac{\widetilde{u}(\mu)-\widetilde{u}(\lambda)}{\mu-\lambda}e^{ix\mu}d\mu\longrightarrow 0\text{ as }x\rightarrow-\infty.$
Thus we have
$-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)e^{ix\mu}}{\mu-\lambda}d\mu\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$
Now we analyze $\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$. According to (2.1), we have
$m_{-}(x,\mu)=\mathrm{e}^{i\mu x}+\mathrm{e}^{i\mu x}K_{\mu}a(x):=\mathrm{e}^{i\mu x}-i\mathrm{e}^{i\mu x}\int_{-\infty}^{x}\mathrm{e}^{-i\mu y}u(y)\Pi\left(\bar{u}\mathrm{e}^{i\mu\cdot}a\right)(y)dy.$
Then
$\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$ $:=-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)e^{ix\mu}}{\mu-\lambda}d\mu-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)\mathrm{e}^{i\mu x}K_{\mu}a(x)}{\mu-\lambda}d\mu$
$:=I_{1}(x,\lambda)+I_{2}(x,\lambda).$
with
$I_{2}(x,\lambda)$
$=-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{2\lambda}\frac{|\widetilde{u}(\lambda)|^{2}\mathrm{e}^{i\mu x}K_{\lambda}a(x)}{\mu-\lambda}d\mu-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{0}^{2\lambda}\frac{\bar{\widetilde{u}}(\lambda)\mathrm{e}^{i\mu x}\left(\widetilde{u}(\mu)K_{\mu}a(x)-\widetilde{u}(\lambda)K_{\lambda}a(x)\right)}{\mu-\lambda}d\mu$
$-\frac{\mathrm{e}^{-i\lambda x}}{4\pi^{2}}p.v.\int_{2\lambda}^{\infty}\frac{\widetilde{u}(\mu)\bar{\widetilde{u}}(\lambda)\mathrm{e}^{i\mu x}K_{\mu}a(x)}{\mu-\lambda}d\mu$
$:=I_{3}(x,\lambda)+I_{4}(x,\lambda)+I_{5}(x,\lambda).$
We already know that
$I_{1}(x,\lambda)\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$ (2.6)
Since $K_{\lambda}a(x)\rightarrow 0$ as $x\rightarrow$-$\infty$, it is not hard to see that
$I_{3}(x,\lambda)\longrightarrow 0\text{ as }x\rightarrow-\infty.$ (2.7)
Since $u\in\mathcal{S}_{+}(\mathbb{R})$ and $\widetilde{u}\in\mathcal{S}_{*}$, we can easily deduce that for $\mu,\lambda>0$,
$|\widetilde{u}(\mu)K_{\mu}a(x)-\widetilde{u}(\lambda)K_{\lambda}a(x)|\leq C(x)|\mu-\lambda|,$
where $C(x)$ does not depend on $\mu,\lambda$ and $C(x)\rightarrow 0$ as $x\rightarrow$-$\infty$.
Then we infer
$|I_{4}|\leq\frac{2\lambda|\widetilde{u}(\lambda)|}{4\pi^{2}}C(x)\rightarrow 0\text{ as }x\rightarrow-\infty.$ (2.8)
Since $\|K_{\mu}a(x)\|_{L_{\mu}^{\infty}}\rightarrow 0$ as $x\rightarrow$-$\infty$ by the Cauchy–Schwarz inequality, we can easily deduce that
$|I_{5}|\leq C_{\lambda}\|\widetilde{u}\|_{L_{\mu}^{2}}\|K_{\mu}a(x)\|_{L_{\mu}^{\infty}}\rightarrow 0\text{ as }x\rightarrow$-$\infty$ (2.9)
Combining (2.6), (2.7), (2.8) and (2.9), we conclude that
$\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)\longrightarrow\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}\text{ as }x\rightarrow-\infty.$
The proof is complete. ∎
From Lemma 2.17, we can then easily give the form of $n(x,\lambda)$.
Corollary 2.18.
Let $u\in\mathcal{S}_{+}(\mathbb{R})$, then we have the following expression of $n(x,\lambda)$
$n(x,\lambda)=n_{1}(x,\lambda)-\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}m_{-}(x,\lambda)$ (2.10)
Proof.
Recall the expression (2.5)
$n(x,\lambda)=n_{1}(x,\lambda)+c\,m_{-}(x,\lambda),$
since $n(x,\lambda)\rightarrow 0$ as $x\rightarrow$-$\infty$, from Lemma 2.17, we can infer that
$c=-\frac{i}{4\pi}|\widetilde{u}(\lambda)|^{2}m_{-}(x,\lambda).$
We keep the assumption $u\in\mathcal{S}_{+}(\mathbb{R})$. Let $\widetilde{f}\in H^{1}(0,\infty)$ and $f\in H_{+}^{1}(\mathbb{R})$, then
[…]
$=-i\widetilde{f}(0_{+})-i\lim_{\lambda\to 0_{+}}\int_{0}^{\infty}\widetilde{f}(\lambda)\,\partial_{\lambda}\chi_{\varepsilon}(\lambda)m_{-}(x,\lambda)d\lambda-i\lim_{\lambda\to 0_{+}}\int_{0}^{\infty}\widetilde{f}(\lambda)\,\chi_{\varepsilon}(\lambda)\partial_{\lambda}m_{-}(x,\lambda)d\lambda$
$=-i\widetilde{f}(0_{+})-i\int_{0}^{\infty}\widetilde{f}(\lambda)\,\partial_{\lambda}m_{-}(x,\lambda)d\lambda$
$=-i\widetilde{f}(0_{+})+x\int_{0}^{\infty}\widetilde{f}(\lambda)\,m_{-}(x,\lambda)d\lambda-i\int_{0}^{\infty}\widetilde{f}(\lambda)n(x,\lambda)d\lambda.$
Here $\chi_{\varepsilon}(\lambda)=\mathrm{e}^{-\varepsilon\lambda^{2}}$, and the second identity comes from the integration by parts and Lemma 2.7.
From Corollary 2.18, we deduce
[…]
$=-\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)n_{1}(x,\lambda)d\lambda-\frac{1}{8\pi^{2}}\int_{0}^{\infty}\widetilde{f}(\lambda)|\widetilde{u}(\lambda)|^{2}m_{-}(x,\lambda)d\lambda$
$=-\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)n_{1}(x,\lambda)d\lambda-\frac{1}{4\pi}\mathcal{F}_{d}^{-1}(\widetilde{f}(\lambda)|\widetilde{u}(\lambda)|^{2}).$
Let us denote the third term above by $r_{3}(x)$. Then we can easily compute its distorted Fourier transform:
$\widetilde{r}_{3}(\mu)=-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)\widetilde{n_{1}}(\mu,\lambda)d\lambda=\frac{i}{4\pi}\widetilde{u}(\mu)H(\bar{\widetilde{u}}\widetilde{f})(\mu).$
Here $H$ is the Hilbert transform on the positive half-line.
Since $u\in\mathcal{S}_{+}(\mathbb{R})$, then we can observe that
$-\frac{i}{2\pi}\int_{0}^{\infty}\widetilde{f}(\lambda)n_{1}(x,\lambda)d\lambda-\frac{1}{4\pi}\mathcal{F}_{d}^{-1}(\widetilde{f}(\lambda)|\widetilde{u}(\lambda)|^{2})=\frac{1}{4\pi}\mathcal{F}_{d}^{-1}\left(i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{f})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{f}(\lambda)\right)\in L_{+}^{2}(\mathbb{R}).$
Also we have $\mathcal{F}_{d}^{-1}(\partial_{\lambda}\widetilde{f})\in L_{+}^{2}(\mathbb{R})$, therefore
$-\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf\in L_{+}^{2}(\mathbb{R}).$
Then from the definition of $X^{*}$ in (1.5) and the identity (1.7), we can deduce that $f\in\operatorname{Dom}(X^{*})$, $\frac{i}{2i\pi}\widetilde{f}(0_{+})+xf=X^{*}f$ and
$\widetilde{f}(0_{+})=\widehat{f}(0_{+}).$ (2.11)
Thus we have
$i\partial_{\lambda}\widetilde{f}=\mathcal{F}_{d}\left(X^{*}f\right)+\frac{1}{4\pi}\left(i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{f})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{f}(\lambda)\right).$ (2.12)
Even in the distorted Fourier transform, the resolvent of $X^{*}f+2tL_{u_{0}}$ is difficult to compute. However, we can easily compute the resolvent of $i\partial_{\lambda}+2t\lambda$ with the following form
$\widetilde{h}(t,\lambda):=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}(\eta)d\eta.$ (2.13)
For every $z\in\mathbb{C}_{+}$, it is not hard to see that $\widetilde{h}\in H^{1}(0,\infty)$ and $h\in H_{+}^{1}(\mathbb{R})$ using the Schur test, so we can apply (2.12) for $h$. Now we justify $I_{+}(h)$. From (2.11), we have
$\frac{1}{2i\pi}I_{+}(h)=\widehat{h}(0_{+})=\widetilde{h}(0_{+})=\frac{1}{2\pi}\int_{0}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}(\eta)d\eta=\mathrm{e}^{it\partial_{x}^{2}}v(z),$ (2.14)
where $\widehat{v}(\xi)=\widetilde{u}(\xi)$ for $\xi>0$ and $\widehat{v}(\xi)$ for $\xi<0$.
In fact, we can easily see that the scattering term in Theorem 1.9 is exacly $\mathrm{e}^{it\partial_{x}^{2}}v$ obtained in (2.14) with $u=u_{0}$. We remark that, from the density of $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{2}(\mathbb{R})$, we can obtain (2.13) for $u\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$.
2.4 Extension of the formula
Let $u\in\mathcal{S}_{+}(\mathbb{R})$. From the previous section, we know that for $\widetilde{f}\in H^{1}(0,\infty)$ and $f\in H_{+}^{1}(\mathbb{R})$, we have $f\in\operatorname{Dom}(X^{*})$ with
$I_{+}(f)=J_{+}(\widetilde{f}):=\widetilde{f}(0_{+})$ (2.15)
and
$\mathcal{F}_{d}\left(X^{*}f\right)=i\partial_{\lambda}\widetilde{f}-\frac{1}{4\pi}\left(i\widetilde{u}(\lambda)H(\bar{\widetilde{u}}\widetilde{f})(\lambda)-|\widetilde{u}(\lambda)|^{2}\widetilde{f}(\lambda)\right):=i\partial_{\lambda}\widetilde{f}+B(\widetilde{u})\widetilde{f}$. (2.16)
Furthermore, we know that, for every $\widetilde{f}\in L^{2}(0,\infty)$,
$(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{f}=\mathrm{e}^{it\lambda^{2}}(i\partial_{\lambda}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}\widetilde{f}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{f}(\eta)d\eta.$
Using the Schur test, we can easily see that $(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{f}\in L^{2}(0,\infty)$, so $(i\partial_{\lambda}+2t\lambda-z)^{-1}$ is well-defined on $L^{2}(0,\infty)$. It is also not hard to see that $\partial_{\lambda}$ is a dissipative operator, therefore $\partial_{\lambda}$ is a maximally dissipative operator. Since B(u~)B($\widetilde${u}) is a self-adjoint operator, iB(u~)-iB($\widetilde${u}) is a dissipative operator. In fact, we can easily observe that
$(i\partial_{\lambda}+B(\widetilde{u})+2t\lambda-z)^{-1}=\mathrm{e}^{it\lambda^{2}}(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}.$ (2.17)
As $i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}$ is dissipative and bounded on $L^{2}(0,\infty)$, from the Miyadera–Gustafson theorem 1.5, we can infer that $(i\partial_{\lambda}+B(\widetilde{u})+2t\lambda-z)^{-1}$ is well-defined on $L^{2}(0,\infty)$. It is not hard to see that
$\widetilde{g}:=(i\partial_{\lambda}+B(\widetilde{u})+2t\lambda-z)^{-1}\widetilde{u}$
satisfying $\widetilde{g}\in H^{1}(0,\infty)$ and $g\in H_{+}^{1}(\mathbb{R})$, then from (2.15) and (2.16), we have
$J_{+}(\widetilde{g})=I_{+}(g),$
where
$g=\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u.$
Then from Theorem 1.8, we can deduce that, if $u(t)$ is a solution of (1.1) with $u(0)=u_{0}\in\mathcal{S}_{+}(\mathbb{R})$, then for every $z\in\mathbb{C}_{+}$,
$u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}u_{0}\right]=\frac{1}{2i\pi}J_{+}\left[(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z)^{-1}\widetilde{u}_{0}\right].$ (2.18)
Now we compute $u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)$ with $\widehat{v}_{0}(\xi)=\widetilde{u}_{0}(\xi)$ for $\xi>0$ and $\widehat{v}(\xi)$ for $\xi<0$. From (2.14), (2.18) and the second resolvent identity, we can infer that
$u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)=\frac{1}{2i\pi}J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right].$ (2.19)
In fact, it is not hard to see that $\partial_{\lambda}$ is also a maximally disspative operator on $L^{1}(0,\infty)$ with $D(\partial_{\lambda})=W^{1,1}(0,\infty)$. For $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$, we can infer that $u_{0}\in L_{+}^{p}(\mathbb{R})$ with some $1<p<2$. Then from Corollary 2.14, we can deduce that $\widetilde{u_{0}}\in L^{q}$ with some $2<q<\infty$. Then for $f\in W^{1,1}(0,\infty)$, we have
$\left\|\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\mathrm{e}^{it\lambda^{2}}f\right\|_{L^{1}(0,\infty)}\leq C\|\widetilde{u}_{0}\|_{L^{q}}^{2}\|f\|_{L^{r}}\leq C\|\widetilde{u}_{0}\|_{L^{q}}^{2}\|f\|_{L^{1}}^{\frac{1}{r}}\|f^{\prime}\|_{L^{1}}^{1-\frac{1}{r}},$
here $r$ satisfies $1<r<\infty$. It is not hard to see that $i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\mathrm{e}^{it\lambda^{2}}$ is a dissipative operator on $L^{1}(0,\infty)$. Combine the above estimate with Young’s inequality for products, we can deduce that $i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\mathrm{e}^{it\lambda^{2}}$ is infinitesimally small with respect to $\partial_{\lambda}$. Then from the Miyadera–Gustafson theorem 1.5, we can infer that $\partial_{\lambda}-i\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}$ is a maximally dissipative operator and $(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}$ is well-defined on $L^{1}(0,\infty)$. More precisely, we can deduce that, for any $f\in L^{1}(0,\infty)$,
$(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}f\in W^{1,1}(0,\infty)$
with
$\left\|(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}f\right\|_{W^{1,1}(0,\infty)}\leq C\left\|f\right\|_{L^{1}(0,\infty)}.$
Therefore we have
$\left|J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right]\right|$ (2.20)
$=$ $\,\left|J_{+}\left[(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right]\right|$
$\leq$ $\,C\left\|(i\partial_{\lambda}+\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u})\mathrm{e}^{it\lambda^{2}}-z)^{-1}\mathrm{e}^{-it\lambda^{2}}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{W^{1,1}(0,\infty)}$
$\leq$ $\,C\left\|B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{L^{1}(0,\infty)}$
$\leq$ $\,C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{2}\left\|\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{L^{\infty}(0,\infty)}$
$\leq$ $\,C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{3}$
$\leq$ $\,C\left\|u_{0}\right\|_{L^{2}(\mathbb{R})}.$
Then by the density from $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{2}(\mathbb{R})$, we can obtain (2.19) and (2.20) for $u\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$.
3 Long-time behavior
Now we can prove the following scattering result for the the defocusing (CM-DNLS) (1.1).
Theorem 3.1.
Let $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R}):=\{w\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with some $\alpha>0$. Then the solution $u\in C(\mathbb{R};L_{+}^{2}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
$\lim_{t\rightarrow+\infty}\left\|u(t)-\mathrm{e}^{i\partial_{x}^{2}}v_{0}\right\|_{L^{2}(\mathbb{R})}=0.$ (3.1)
Here $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$, and
$\mathcal{F}_{d}(f)=\widetilde{f},\qquad\forall f\in L_{+}^{2}(\mathbb{R})$
is the distorted Fourier transform assciated with $L_{u_{0}}$.
Proof.
First of all, the mass conservation law $\|u(t)\|_{L^{2}(\mathbb{R})}=\|u_{0}\|_{L^{2}(\mathbb{R})}$ implies that u(t){u(t)} is uniformly bounded in $L_{+}^{2}(\mathbb{R})$, hence it admits weakly convergent subsequences in $L_{+}^{2}(\mathbb{R})$ as $t\to\infty$. We shall prove that the weak limit is unique.
Let us recall the formula (2.19) for $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ with some $\alpha>0$,
$u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)=\frac{1}{2i\pi}J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right].$ (3.2)
We recall (2.13) for $u=u_{0}$ and we rewrite it as
$\widetilde{h}_{0}(t,\lambda):=(i\partial_{\lambda}+2t\lambda-z)^{-1}\widetilde{u}_{0}=i\,\mathrm{e}^{it\lambda^{2}-iz\lambda}\int_{\lambda}^{\infty}\mathrm{e}^{-it\eta^{2}+iz\eta}\widetilde{u}_{0}(\eta)d\eta.$
Then from (2.20) and (3.2), we infer
$\left|u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)\right|$
$\leq\,$ $C\left|J_{+}\left[\left(i\partial_{\lambda}+B(\widetilde{u}_{0})+2t\lambda-z\right)^{-1}B(\widetilde{u}_{0})\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right]\right|$
$\leq\,$ $C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{2}\left\|\left(i\partial_{\lambda}+2t\lambda-z\right)^{-1}\widetilde{u}_{0}\right\|_{L^{\infty}(0,\infty)}$
$\leq\,$ $C\left\|\widetilde{u}_{0}\right\|_{L^{2}(0,\infty)}^{2}\|\widetilde{h}_{0}(t)\|_{L^{\infty}(0,\infty)}.$
Since $z\in\mathbb{C}_{+}$, Now we apply the Riemann–Lebesgue lemma and we can conclude that
$\lim_{t\to\infty}\|\widetilde{h}_{0}(t)\|_{L_{\lambda}^{\infty}(0,\infty)}=0$
Thus we deduce
$\forall z\in\mathbb{C}_{+},\qquad u(t,z)-\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z)\to 0\text{ as }t\to\infty.$
If $u(t_{n})\rightharpoonup w$ in $L_{+}^{2}(\mathbb{R})$ as $t_{n}\to\infty$, then for every $z\in\mathbb{C}_{+}$, we have
$u(t_{n},z)-w(z)\to 0\text{ as }t_{n}\to\infty.$
Therefore we can deduce $w=\mathrm{e}^{it\partial_{x}^{2}}v_{0}$ and the weak limit is unique, i.e.,
$u(t)\rightharpoonup\mathrm{e}^{it\partial_{x}^{2}}v_{0}\text{ in }L_{+}^{2}(\mathbb{R})\text{ as }t\to\infty.$
Since $\|u(t)\|_{L^{2}}=\|u_{0}\|_{L^{2}}=\|\mathrm{e}^{it\partial_{x}^{2}}v_{0}\|_{L^{2}}$, we conclude that this weak convergence is exactly the strong convergence in $L_{+}^{2}(\mathbb{R})$. The proof is complete. ∎
Remark 3.2.
Following the approach developed in Section 2.3 and Section 3, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions with initial data in $L_{+}^{2,\alpha}(\mathbb{R}):=\{u\in L_{+}^{2}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for some $\alpha>0$.
We fix $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})$ for some $\alpha>0$. From [12, Section 5], we know that every eigenvalue of $L_{u_{0}}=D-u_{0}T_{\bar{u}_{0}}$ (the Lax operator of (1.3)) is simple, and the number $N$ of eigenvalues of $L_{u_{0}}$ is finite with $N\leq\frac{\|u_{0}\|_{L^{2}}^{2}}{2\pi}$. From [11, Theorem 4.3], we know that for every $f\in L_{+}^{2}(\mathbb{R})$
$f(x)=\sum_{j=1}^{N}\left\langle f,\varphi_{j}\right\rangle\varphi_{j}(x)+\mathcal{F}_{d}^{-1}(\widetilde{f})(x)=P_{p}(L_{u_{0}})f(x)+P_{ac}(L_{u_{0}})f(x).$ (3.3)
Here $\left\{\varphi_{j}\right\}_{j=1}^{N}$ are the normalized eigenfunctions of $L_{u_{0}}$, $\mathcal{F}_{d}(f)=\widetilde{f}$ is the distorted Fourier transform defined in [11, (4.1)], $P_{p}(L_{u_{0}})$ denote the projection onto the eigensubspace of $L_{u_{0}}$ and $P_{ac}(L_{u_{0}})$ denote the projection onto the absolutely continuous subspace of $L_{u_{0}}$. Now we apply (3.3) for $f=u_{0}$, and we obtain
$u_{0}(x)=P_{p}(L_{u_{0}})u_{0}(x)+P_{ac}(L_{u_{0}})u_{0}(x).$ (3.4)
Now we assume that the solution $u(t,\cdot)\in L_{+}^{2}(\mathbb{R})$ to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}$ exists globally in time (this is always true if $\|u_{0}\|_{L^{2}}^{2}<2\pi$ or if $u_{0}\in H_{+}^{1}(\mathbb{R})$ with $\|u_{0}\|_{L^{2}}^{2}=2\pi$). From the explicit formula (3.2), we infer that for every $t\in\mathbb{R}$ and every $z\in\mathbb{C}_{+}$,
$u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}\varphi\right]+\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}P_{ac}(L_{u_{0}})u_{0}\right].$ (3.5)
For the second term above, following the approach in the proof of Theorem 3.1, we can show that
$\forall z\in\mathbb{C}_{+},\qquad\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}P_{ac}(L_{u_{0}})u_{0}\right]\rightarrow\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z),$
where $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$. From this we can deduce the sacttering term of the focusing (CM-DNLS) (1.3). Notice that if $\|u_{0}\|_{L^{2}}^{2}<2\pi$, $L_{u_{0}}$ has no eigenvalues. In this case $P_{p}(L_{u_{0}})=0$ and $P_{ac}(L_{u_{0}})u_{0}=u_{0}$, so from the above argument we can directly deduce the scattering of the focusing (CM-DNLS) (1.3) for $u_{0}\in L_{+}^{2,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ with some $\alpha>0$.
Now we analyse the first term on the right hand side in (3.5). Following the method used in the proof of [9, Lemma 3.1], we can also deduce that
$\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}P_{p}(L_{u_{0}})u_{0}\right]=U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}P_{p}(L_{u_{0}})u_{0}(z).$
and
$u(t)=U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}u_{0}.$
Since $\left\{\varphi_{j}\right\}_{j=1}^{N}$ are the normalized eigenfunctions of $L_{u_{0}}$, we know that
$U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}P_{p}(L_{u_{0}})u_{0}$
$=$ $\,\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}\varphi_{j}$
$=$ $\,\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\mathrm{e}^{-it\lambda_{j}^{2}}U(t)\varphi_{j}.$
Here the unitary operator $U(t)$ is defined by the linear initial value problem in $\mathscr{L}\left(L_{+}^{2}(\mathbb{R})\right)$,
$U^{\prime}(t)=B_{u(t)}U(t),\qquad U(0)=\operatorname{Id}\text{ with }B_{u}:=uT_{\partial_{x}\bar{u}}-\partial_{x}uT_{\bar{u}}+i(uT_{\bar{u}})^{2}.$
From the Lax identity $L_{u(t)}U(t)=U(t)L_{u_{0}}$, we can infer that $\phi_{j}(t):=\mathrm{e}^{-it\lambda_{j}^{2}}U(t)\varphi_{j}$ is the normalized eigenfunctions of $L_{u(t)}$. Thus we have
$U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}P_{p}(L_{u_{0}})u_{0}=\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\phi_{j}(t).$
So we have
$u(t)=\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\phi_{j}(t)+\mathrm{e}^{it\partial_{x}^{2}}v_{0}+o(1)$ (3.6)
in $L_{+}^{2}(\mathbb{R})$. Notice that (3.6) is exactly the asymptotic bound-state/radiation decomposition for the global solution $u(t)$ to the focusing (CM-DNLS) (1.3). In fact, we have
$\left\langle u_{0},\varphi_{j}\right\rangle=\left\langle U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}u_{0},U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}\varphi_{j}\right\rangle=\left\langle u(t),\phi_{j}(t)\right\rangle,$
therefore $\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\phi_{j}(t)=P_{p}(L_{u(t)})(u(t))$. Moreover, from [11, Lemma 3.7], we know that for every $1\leq j\leq N$,
$|\left\langle u_{0},\varphi_{j}\right\rangle|^{2}=2\pi.$
Then we can deduce that
$\|v_{0}\|_{L^{2}}^{2}=\|P_{ac}(L_{u_{0}})u_{0}\|_{L^{2}}^{2}=\|u_{0}\|_{L^{2}}^{2}-\sum_{j=1}^{N}|\left\langle u_{0},\varphi_{j}\right\rangle|^{2}=\|u_{0}\|_{L^{2}}^{2}-2\pi N,$
which coincides with one of the statements in [21, Theorem 1.1].
Remark $3.$
From the soliton resolution result in [21, Theorem 1.1] and Remark 3.2, we are naturally led to the following conjecture: as $t\to\infty$, $\sqrt{2\pi}\phi_{j}(t)$ should asymptotically converge to a traveling soliton profile.